### v1

# Meta-automatic Sequences

Abstract
Nested (or meta-Fibonacci) recurrences, such as the recurrence used to define Hofstadter’s $Q$-sequence, along with the digit-based recurrences that underlie automatic sequences are of interest from both number-theoretic and combinatorial points of view. In this direction, Allouche and Shallit showed how the frequency sequence of a variant of the $Q$-sequence is $2$-automatic. This inspires us to introduce what may be seen as a natural combination of the recurrences for meta-Fibonacci and automatic sequences, by introducing the concept of a meta-automatic sequence. We show how it is possible to construct a meta-automatic sequence that is not denestable in terms of not being reducible, in a specific sense formalized in this paper, to certain digit-based recurrences for automatic sequences. This motivates our study of the non-denestable, meta-automatic sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ introduced in this paper. For each of these integer sequences, we prove explicit DFAO evaluations, together with $4$-uniform morphisms, and we also consider the factor complexities of these sequences.
1 Introduction
Informally, a sequence $(a(n))_{n\geq 0}$ is said to satisfy a nested or meta-Fibonacci recurrence if $a(n)$, for all sufficiently large $n$, can be expressed in terms of $a$ evaluated at an expression involving a value given as a term in $(a(n))_{n\geq 0}$. The classical and prototypical example is Hofstadter’s $Q$-sequence [19, p. 137], which is indexed in the On-Line Encyclopedia of Integer Sequences [28] as A005185, and which satisfies $Q(1)=Q(2)=1$ and
Q(n)=Q(nQ(n1))+Q(nQ(n2))Q(n)=Q(n-Q(n-1))+Q(n-Q(n-2)) (1)
for $n>2$. In contrast to nested recurrences, digit-based recurrences (which may also be referred to as divide-and-conquer recurrences) define $a\big(k^{t}n+j\big)$ (for $j\in\{0,1,\ldots,k^{t}-1\}$) as a function of previously computed values of the form $a\big(k^{t^{\prime}}n^{\prime}+j^{\prime}\big)$ for $t^{\prime}<t$ (and for $j^{\prime}\in\{0,1,\ldots,k^{t^{\prime}}-1\}$). These digit-based recurrences are used to define automatic sequences [3, 9], which are of great importance in many areas of number theory, computer science, and combinatorics.
The depth of the study of meta-Fibonacci sequences is evidenced by the many open problems related to Hofstadter’s $Q$-sequence and generalizations. Notably, to this day, it is not even known whether or not $Q(n)$ is defined for all integers $n>0$. This together with the roles played by automatic sequences within many disciplines in mathematics and computer science lead us to consider a new and interdisciplinary subject that may be thought of as “combining” the recursions for meta-Fibonacci and automatic sequences.
As it turns out, relationships between meta-Fibonacci and automatic sequences have previously been considered. Notably, Allouche and Shallit [4], in 2012, showed that the frequency sequence of a variant of Hofstadter’s $Q$-sequence satisfies a digit-based recurrence, so that this variant sequence is $2$-automatic. In this context, they proved an automaticity criterion [4, Theorem 2.2] reviewed in Section 1.2 below.
In this paper, we introduce the concept of a meta-automatic sequence, which, informally, can be thought of as extending both the Allouche–Shallit criterion and the recursions for meta-Fibonacci sequences as in (1). For example, in contrast to a rule such as $a(4n+2)=a(2n)$ used to define an automatic sequence, we consider rules such as
a(4n+2)=a(2n+1a(n)),a(4n+2)=a(2n+1-a(n)), (2)
where the argument of the right-hand side depends on the value $a(n)$. Recurrences as in (2) do not, in general, fall under the Allouche–Shallit automaticity criterion [4, Theorem 2.2], since the argument on the right-hand side of (2) is not necessarily equivalent to a polynomial in $n$ of the desired form.
The idea of combining the recursions associated with automatic and meta-Fibonacci sequences is motivated by many past references concerning meta-Fibonacci sequences, in addition to Allouche and Shallit’s work on the automaticity of the frequency sequence for Hofstadter’s $Q$-sequence [4]. In this direction, we highlight notable references on the Hofstadter $Q$-sequence [16, 17, 18, 25], and on variants and generalizations of the Hofstadter $Q$-sequence [1, 2, 5, 12, 14, 30] and on meta-Fibonacci sequences [7, 8, 10, 11, 13, 15, 20, 21, 22, 26, 29].
1.1 Outline
A formal definition of the term meta-automatic sequence is given in Section 2 below.
As we demonstrate, it is possible for the recurrence for a meta-automatic sequence to, in a specific sense, degenerate in such a way so that it reduces to a digit-based recurrence. To formalize what should be meant by such reductions, we introduce the concept of denestability. As it turns out, there exist meta-automatic sequences that are not denestable, and we introduce the first construction of a sequence of this form. In this direction, Theorem 23 shows that the integer sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ that we introduce in this paper are not denestable. This leads to the following question motivating the topic of our paper: Under what conditions do meta-automatic recurrences produce automatic sequences?
The binary sequences $(a(n))_{n\geq 0}$ under consideration in this paper are mainly balanced in the sense that $a(2n)+a(2n+1)=1$. This condition provides a defining property of the Thue–Morse sequence, and so the meta-automatic sequences introduced in this paper may be seen as “meta-Thue–Morse” sequences.
We prove that in base $2$, meta-automatic sequences that are binary and balanced are ultimately periodic (Lemma 14). However, as we later clarify, by passing to a base-$4$ digit template that preserves a dyadic balance, this can be used to construct balanced, binary, and meta-automatic sequences that are not ultimately periodic. In this direction, a key idea is given by how the balance constraint linearizes each meta-automatic selector via an $\mathbb{F}_{2}$-affine operation (Lemma 10). This motivates our study of the following meta-automatic sequences:
  • Meta(1) $\mathcal{M}_{1}$ (A392736): a mixed template in which one of the two branch rules uses a meta-automatic selector. The XOR linearization yields a $4$-state DFAO (Section 3).
  • Meta(2) $\mathcal{M}_{2}$ (A391614): a “pure” two-selector template in which both branch rules are meta-automatic. The linearized recurrence has an affine structure over $\mathbb{F}_{2}^{2}$ governed by the matrix $\bigl(\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\bigr)$, and the sequence admits the factorization $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, where $q$ is an explicit bit-masking operator and $\mathbf{t}$ denotes the Thue–Morse sequence (Theorem 35).
Remarkably, despite $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ sharing the same automaton size of $4$ states, the two sequences differ sharply in structural complexity. Notably, we find that $\mathcal{M}_{2}$ admits the evaluation $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, while $\mathcal{M}_{1}$ requires a lifted size-two kernel. Despite this shared automaton size, their factor-complexity functions, studied in Section 5, differ markedly.
For each of the above sequences, we give an MSB-first, base-$4$ DFAO, an explicit $4$-uniform morphism, and factor-complexity results. The broader theme is that the interplay between balancedness and $\mathbb{F}_{2}$-linearity can bring meta-automatic, digit-based recurrences into the realm of automatic sequences, even when criteria for automaticity, such as the above referenced Allouche–Shallit criterion [4], do not apply.
1.2 Preliminaries
For background on automatic sequences, we refer to the appropriate texts from Allouche and Shallit [3] and from Shallit [27]. We briefly recall some definitions that are useful or neeed, for our purposes. In this direction, the concept of a $k$-kernel [3, §6.6] is required in our work.
Definition 1.
Let $U=(U(n))_{n\geq 0}$ denote an infinite sequence. The $k$-kernel of $(U(n))_{n\geq 0}$ [3, p. 185] is
$K_{k}(U)=\left\{\left(U_{k^{i}n+j}\right)_{n\geq 0}:\text{$i\geq 0$ and $0\leq j<k^{i}$}\right\}.$
For $k\geq 2$, a $k$-automatic sequence may then be defined as a sequence $U=(U(n))_{n\geq 0}$ such that $K_{k}(U)$ is finite. Equivalently, this sequence is $k$-automatic if there exists a DFAO that, on input the base-$k$ representation of $n$, outputs $U(n)$, again referring to Allouche and Shallit’s text for background.
Definition 2.
Fix a base $k\geq 2$. By a digit-based recurrence we mean a system of recurrences that defines $a(kn+r)$ for digits $r\in\{0,1,\ldots,k-1\}$ in terms of values of $a(m)$ at strictly smaller indices $m<kn+r$ (typically among $m=n,2n,\dots$).
Let $\mathbf{t}:\mathbb{N}_{0}\to\{0,1\}$ denote the Thue–Morse sequence, defined by $\mathbf{t}(0)=0$ and
$\mathbf{t}(2n)=\mathbf{t}(n),\qquad\mathbf{t}(2n+1)=1-\mathbf{t}(n)\qquad(n\geq 0).$
Equivalently, $\mathbf{t}(n)=s_{2}(n)\bmod 2$, where $s_{2}(n)$ is the sum of the binary digits of $n$.
Since much of our work is closely related to an automaticity criterion due to Allouche and Shallit [4], it is necessary, for our purposes, to review this criterion, as below.
Let $U=(U(n))_{n\geq 0}$ denote a sequence, and let $\alpha$ denote an integer. Adopting notation from Allouche and Shallit’s work [4], for $n\geq-\alpha$, we write $U^{\alpha}$ in place of the sequence given by
$U^{\alpha}:=U(n+\alpha).$
Similarly, letting $q$ and $i$ and $j$ be positive integers, we write $U_{q,i,j}$ in place of the sequence defined by
$U_{q,i,j}(n)=U\left(q^{i}n+j\right)$ (3)
for $n\geq 0$. Moreover, for integers $q\geq 2$ and $t\geq 1$, we write $s(q,t)=\frac{1-q^{t+1}}{1-q}$ in place of the number of pairs $(i,j)$ of integers $i$ and $j$ such that $0\leq i\leq t$ and $0\leq j\leq q^{i}-1$. We then order expressions of the form $U_{q,i,j}$ by writing $U_{1}=U$, $U_{2}$, $\ldots$, $U_{s(q,t)}$. We also let $m(q,t)=q^{t+1}-1$.
Theorem 3.
(Allouche &\& Shallit, 2012) Let the terms of $(U(n))_{n\geq 0}$ be in a finite set $\mathcal{A}$, and let $q\geq 2$ denote an integer. Then $U$ is $q$-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ together with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions from $\mathcal{A}^{a+b+s(q,t)}$ to $\mathcal{A}$ such that
$U\big(q^{t+1}n+j\big)=f_{j}\big(U^{-a}(n),\ldots,U^{-1}(n),U^{0}(n),U^{1}(n),\ldots,\\ U^{b}(n),U_{2}(n),U_{3}(n),\ldots,U_{s(q,t)}(n)\big)$
for all $n\geq n_{0}$ and for all $j\in\{0,1,\ldots,m(q,t)\}$ [4].
2 Meta-automatic sequences
Let $q>0$ and $i_{0}>0$ and $j_{0}\geq 0$ denote integers. For some $v\geq 0$, and for indices $k\in\{1,2,\ldots,v\}$, let $\mathcal{S}_{k}$ and $i_{k}>0$ and $j_{k}\geq 0$ be integers, subject to the condition that $i_{k}<i_{0}$ for all indices $k$. By analogy with (3), and with the assumption that the codomain of $U$ is contained in $\mathbb{N}_{0}$, we write
$U_{q,i_{0},j_{0},\mathcal{S}_{1},i_{1},j_{1},\ldots,\mathcal{S}_{v},i_{v},j_{v}}=U\Big(q^{i_{0}}n_{0}+j_{0}+\mathcal{S}_{1}U(q^{i_{1}}n_{1}+j_{1})+\cdots+\mathcal{S}_{v}U(q^{i_{v}}n_{v}+j_{v})\Big).$
Definition 4.
Let the terms of $(U(n))_{n\geq 0}$ be in a finite subset $\mathcal{A}$ of $\mathbb{N}_{0}$. Let $q\geq 2$ be an integer. Then $U$ is $q$-meta-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ along with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions to $\mathcal{A}$ (with domains given below) such that, for all $n\geq n_{0}$ and for all $j\in\{0,1,\ldots,m(q,t)\}$, the value $U\big(q^{t+1}n+j\big)$ is equal to $f_{j}$ evaluated at a tuple of the form
$\Big(U^{-a}(n),\ldots,U^{-1}(n),U^{0}(n),U^{1}(n),\ldots,U^{b}(n),\\ U_{q,i_{0}^{(1)},j_{0}^{(1)},\mathcal{S}_{1}^{(1)},i_{1}^{(1)},j_{1}^{(1)},\ldots,\mathcal{S}_{v_{1}}^{(1)},i_{v_{1}}^{(1)},j_{v_{1}}^{(1)}},\ldots,U_{q,i_{0}^{(w)},j_{0}^{(w)},\mathcal{S}_{1}^{(w)},i_{1}^{(w)},j_{1}^{(w)},\ldots,\mathcal{S}_{v_{w}}^{(w)},i_{v_{w}}^{(w)},j_{v_{w}}^{(w)}}\Big)$
for some $w\geq 0$ and for integers $i_{k_{1}}^{(k_{2})}\leq t$ and $j_{k_{1}}^{(k_{2})}$ and $\mathcal{S}_{k_{1}}^{(k_{2})}$, and where the domain of $f_{j}$ is $\mathcal{A}^{a+b+w+1}$.
Example 5.
Let the integer sequence $(\mathcal{T}(n))_{n\geq 0}$ be defined so that $\mathcal{T}(0)=0$ and so that
$\mathcal{T}(2n)$ $=\mathcal{T}(n-\mathcal{T}(n-1))\ \text{for $n\geq 1$, and}$
$\mathcal{T}(2n+1)$ $=1-\mathcal{T}(n)\ \text{for $n\geq 0$}.$
According to Definition 4, the sequence $(\mathcal{T}(n))_{n\geq 0}$ is 22-meta-automatic. It can be shown that this sequence agrees, for $n\geq 1$, with the OEIS sequence A039982, which is defined in a different way and without the use of meta-recursions.
Definition 6.
For a $q$-meta-automatic sequence $(U(n))_{n\geq 0}$, adopting notation from Definition 4, a selector refers to an expression given by $f_{j}$ evaluated at the tuple displayed in Definition 4, i.e., with such an expression being equal to $U(q^{t+1}n+j)$.
Example 7.
For the $2$-meta-automatic sequence in Example 5, the expression $\mathcal{T}(n-\mathcal{T}(n-1))$ is a selector.
Binary sequences of the following form are a main object of study in our work.
Definition 8.
A binary sequence $a:\mathbb{N}_{0}\to\{0,1\}$ is said to be balanced if
$a(2n)+a(2n+1)=1\qquad(n\geq 0).$ (4)
Example 9.
The Thue–Morse sequence may be seen as a prototypical instance of a balanced (binary) sequence.
As in the following result, we let $\oplus$ denote the additive operation in the field $\mathbb{F}_{2}$.
Lemma 10.
If $a$ is balanced, then
$a(2n+a(n))=a(2n)\oplus a(n)\quad\text{and}\quad a(2n+1-a(n))=a(2n+1)\oplus a(n),$
for all $n\geq 0$.
Proof.
This follows in a direct way from (4). ∎
Remark 11.
The selector a(2n+1a(n))a(2n+1-a(n)) and the XOR expression $a(2n+1)\oplus a(n)$ compute the same function of the triple $(a(n),a(2n),a(2n+1))$. Both read the same two independent bits $(a(n),a(2n))$, since balancedness gives a(2n+1)=1a(2n)a(2n+1)=1-a(2n). The balance constraint does not reduce the number of value dependencies, but it ensures that the dependency is $\mathbb{F}_{2}$-affine: The selector, viewed as a Boolean function, is addition in $\mathbb{F}_{2}$.
2.1 Denestability
We introduce two notions that distinguish the sequences studied in this paper from those covered by the automaticity criterion of Allouche and Shallit [4, Theorem 2.2].
Definition 12.
A meta-automatic recurrence for a sequence $(U(n))_{n\geq 0}$ is denestable if every selector associated with an expression of the form $U\big(q^{t+1}n+j\big)$ can, for sufficiently large $n\geq n_{0}$, be written as an expression of the form given in the right-hand side of the identity in Theorem 3, without increasing the size of kernel.
Example 13.
The integer sequence $(\mathcal{T}(n))_{n\geq 0}$ defined in Example 5 is denestable, since the given recursion is such that $\mathcal{T}(2n)=1$ for all $n\geq 1$, providing, together with the relation $\mathcal{T}(2n+1)=1-\mathcal{T}(n)$, a digit-based recurrence for $\mathcal{T}$.
A weaker notion, relative to Definition 12, is that of $\mathbb{F}_{2}$-linearizability, which, informally, allows the balance constraint to convert meta-automatic selectors into $\mathbb{F}_{2}$-affine maps at fixed arguments. This is later clarified.
In base $2$, the most direct way to introduce a meta-automatic argument while keeping the binary balancedness (4) is, informally, to try to define the even subsequence using a “one-level” selector at the $n$-scale. For instance, set
$a(2n)=a(n+a(n))\qquad\text{or}\qquad a(2n)=a(n+1-a(n)),$
and then set a(2n+1)=1a(2n)a(2n+1)=1-a(2n) to preserve (4). Under balance, however, both schemes collapse to ultimate periodicity, and this is formalized below.
Lemma 14.
Let $a:\mathbb{N}_{0}\to\{0,1\}$ be balanced.
  1. (a)
    If $a(2n)=a(n+a(n))$ for all $n\geq 0$, then $a(0)=0$ and
    $a(2n)=0,\qquad a(2n+1)=1\qquad(n\geq 2),$
    so $a$ is ultimately periodic of period $2$.
  2. (b)
    If a(2n)=a(n+1a(n))a(2n)=a(n+1-a(n)) for all $n\geq 0$, then $a(0)=1$ and
    $a(2n)=1,\qquad a(2n+1)=0\qquad(n\geq 2),$
    so $a$ is ultimately periodic of period $2$.
Proof.
(a) Taking $n=0$ gives $a(0)=a(a(0))$. If $a(0)=1$, then balancedness forces $a(1)=0$, and hence $a(0)=a(1)=0$, a contradiction. Thus $a(0)=0$ and $a(1)=1$.
Write $b_{m}=a(2m)$, so that balancedness gives $a(2m+1)=1-b_{m}$. Using $a(2n)=a(n+a(n))$ with $n=2m$, we obtain
$b_{2m}=a(4m)=a(2m+a(2m))=a(2m+b_{m})=\begin{cases}a(2m)=b_{m},&b_{m}=0;\\[2.0pt] a(2m+1)=1-b_{m}=0,&b_{m}=1.\end{cases}$
So, we have that $b_{2m}=0$ for all $m\geq 0$.
Next, using $a(2n)=a(n+a(n))$ with $n=2m+1$ and $a(2m+1)=1-b_{m}$, we get
$b_{2m+1}=a(4m+2)=a(2m+1+a(2m+1))=a(2m+2-b_{m})=\\ \begin{cases}a(2m+2)=b_{m+1},&b_{m}=0;\\[2.0pt] a(2m+1)=0,&b_{m}=1.\end{cases}$
We claim that $b_{m}=0$ for all $m\geq 2$. This holds for $m=2$, since $b_{2}=0$, and for $m=3$ because the above identity with $m=1$ gives $b_{3}=b_{2}$ if $b_{1}=0$, and $b_{3}=0$ if $b_{1}=1$; in either case $b_{3}=0$ since $b_{2}=0$. Now let $m\geq 4$. If $m$ is even then $b_{m}=b_{2r}=0$. If $m$ is odd, write $m=2r+1$ with $r\geq 1$. If $b_{r}=1$ then $b_{m}=0$; if $b_{r}=0$ then $b_{m}=b_{r+1}$, and $r+1<m$, so by induction $b_{r+1}=0$. Hence $b_{m}=0$ for all $m\geq 2$, i.e., $a(2n)=0$ for all $n\geq 2$. The balance condition in (4) then gives $a(2n+1)=1$ for all $n\geq 2$.
(b) This follows from (a) by complementing: if $a$ satisfies (b) then $\bar{a}(n)=1-a(n)$ satisfies (a), and vice versa. ∎
Lemma 14 shows that meta-automatic behavior requires passing to base 44: the dyadic pairing (4) is preserved, but there is now room for selectors one scale below.
2.2 A non-nested baseline case
Fix $a(0)=0$, $a(1)=1$, and define
$a(4n)=F(n),\quad a(4n+1)=1-F(n),a(4n+2)=G(n),\quad a(4n+3)=1-G(n),$ (5)
so that $a(2n)+a(2n+1)=1$ holds automatically. The three main sequences investigated in this paper arise by specializing (5). To begin with, we consider an integer sequence $(\mathcal{Q}(n))_{n\geq 0}$ that does not have a nested structure in terms of its recurrence but that is to be used extensively in our subsequent work.
Definition 15.
The Thue–Morse Quarto sequence $\mathcal{Q}:\mathbb{N}_{0}\to\{0,1\}$ (A298952, where we use the convention $\mathcal{Q}(0)=0$ rather than $a(0)=1$ as in the OEIS entry) is defined by $\mathcal{Q}(0)=0$, $\mathcal{Q}(1)=1$, and
$\mathcal{Q}(4n)$ $=\mathcal{Q}(n),$ $\mathcal{Q}(4n+1)$ $=1-\mathcal{Q}(n),$ (6)
$\mathcal{Q}(4n+2)$ $=\mathcal{Q}(2n),$ $\mathcal{Q}(4n+3)$ $=1-\mathcal{Q}(2n).$ (7)
Omitting details for brevity, we have that there is a $3$-state DFAO for $(\mathcal{Q}(n))_{n\geq 0}$, in such a way so that $\mathcal{Q}$ is $4$-automatic and computed by the DFAO with states $\{A,B,C\}$, initial state $A$, and output $\tau(A)=\tau(B)=0$, $\tau(C)=1$. This DFAO is shown in Figure 1. The DFAO for $\mathcal{Q}$ corresponds to the $4$-uniform morphism
$A\mapsto ACBC,\quad B\mapsto BCCB,\quad C\mapsto CBBC,$
with coding $\tau(A)=\tau(B)=0$, $\tau(C)=1$.
A/0A/0B/0B/0C/1C/100,30,30,30,3221,31,31,21,21,21,2
Figure 1: DFAO for $\mathcal{Q}$ (MSB-first). State labels are state/output. The initial state is $A$.
Proposition 16.
Define $d:\mathbb{N}_{0}\to\{0,1\}$ by $d(0)=0$ and $d(n)=\lfloor\log_{2}n\rfloor\bmod 2$ for $n\geq 1$. Then
$\mathcal{Q}(n)=\mathbf{t}(n)\oplus d(n)\qquad(n\geq 0).$
Proof.
The cases $n=0,1$ are immediate. For $n\geq 1$ and $r\in\{0,1,2,3\}$, since $\lfloor\log_{2}(4n+r)\rfloor=2+\lfloor\log_{2}n\rfloor$, we have $d(4n+r)=d(n)$ for all $r$. We also use $\mathbf{t}(4n+r)=\mathbf{t}(n)\oplus\mathbf{t}(r)$, and we proceed by induction.
Case $r=0$: $\mathcal{Q}(4n)=\mathcal{Q}(n)=\mathbf{t}(n)\oplus d(n)$ and $\mathbf{t}(4n)\oplus d(4n)=\mathbf{t}(n)\oplus d(n)$.
Case $r=1$: $\mathcal{Q}(4n+1)=1\oplus\mathcal{Q}(n)=1\oplus\mathbf{t}(n)\oplus d(n)$ and $\mathbf{t}(4n+1)\oplus d(4n+1)=(1\oplus\mathbf{t}(n))\oplus d(n)$.
Case $r=2$: $\mathcal{Q}(4n+2)=\mathcal{Q}(2n)$. By induction, $\mathcal{Q}(2n)=\mathbf{t}(2n)\oplus d(2n)=\mathbf{t}(n)\oplus(d(n)\oplus 1)=1\oplus\mathbf{t}(n)\oplus d(n)$. Also $\mathbf{t}(4n+2)=\mathbf{t}(2n+1)=1\oplus\mathbf{t}(n)$, so $\mathbf{t}(4n+2)\oplus d(4n+2)=(1\oplus\mathbf{t}(n))\oplus d(n)$.
Case $r=3$: $\mathcal{Q}(4n+3)=1\oplus\mathcal{Q}(2n)=\mathbf{t}(n)\oplus d(n)$, and $\mathbf{t}(4n+3)\oplus d(4n+3)=\mathbf{t}(n)\oplus d(n)$. ∎
3 Meta(1)
Definition 17 (Meta(1)).
Define $\mathcal{M}_{1}:\mathbb{N}_{0}\to\{0,1\}$ by $\mathcal{M}_{1}(0)=0$, $\mathcal{M}_{1}(1)=1$, and for $n\geq 0$, so that
$\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n),\ \mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(n),$ (8)
and
$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),\ \mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (9)
Lemma 18.
The sequence $\mathcal{M}_{1}(n)$ is well-defined for all $n\geq 0$ and satisfies $\mathcal{M}_{1}(2n)+\mathcal{M}_{1}(2n+1)=1$.
Proof.
Well-definedness follows by induction on $m$. For $m=4n+r$, the right-hand sides in (8) reduce to indices <m<m; for (9), the selector index is $2n+1-\mathcal{M}_{1}(n)\in\{2n,2n+1\}<4n+2\leq m$. Balancedness follows from (5), with $\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(4n)$ and $\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(4n+2)$. ∎
Remark 19.
Because $\mathcal{M}_{1}$ is balanced, Lemma 10 yields
$\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n),$
so (9) is equivalent to the non-nested relation $\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$. The selector form is the natural “meta” description and is the form that generalizes.
3.1 A 4-state DFAO
Define the auxiliary function
$b(n):=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (10)
Consider the lifted state $s(n):=(\mathcal{M}_{1}(n),b(n))\in\{0,1\}^{2}$.
Lemma 20.
For each $r\in\{0,1,2,3\}$, the state $s(4n+r)$ depends only on $s(n)$.
Proof.
Write $s(n)=(x,y)$ where $x=\mathcal{M}_{1}(n)$ and $y=b(n)$. By the balancedness given by Lemma 18 and the XOR linearization (Lemma 10), we have $y=b(n)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)=(1-\mathcal{M}_{1}(2n))\oplus x$, so that
$\mathcal{M}_{1}(2n)=1\oplus x\oplus y,\qquad\mathcal{M}_{1}(2n+1)=x\oplus y.$ (11)
Case $r=0$. The first coordinate is $\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n)=x$. For the second coordinate, $b(4n)=\mathcal{M}_{1}(8n+1-\mathcal{M}_{1}(4n))=\mathcal{M}_{1}(8n+1-x)$. If $x=0$, then $b(4n)=\mathcal{M}_{1}(8n+1)=\mathcal{M}_{1}(4(2n)+1)=1-\mathcal{M}_{1}(2n)=1-(1\oplus y)=y$. If $x=1$, then $b(4n)=\mathcal{M}_{1}(8n)=\mathcal{M}_{1}(4(2n))=\mathcal{M}_{1}(2n)=1\oplus x\oplus y=y$. In both cases, we find that s(4n)=(x,y)s(4n)=(x,\,y).
Case $r=1$. The first coordinate is $\mathcal{M}_{1}(4n+1)=1-x=1\oplus x$. For the second coordinate, we have that $b(4n+1)=\mathcal{M}_{1}(8n+2+x)$. If $x=0$, then $b(4n+1)=\mathcal{M}_{1}(8n+2)=\mathcal{M}_{1}(4(2n)+2)=b(2n)$. Using the property that $b(2n)=\mathcal{M}_{1}(4n+1-\mathcal{M}_{1}(2n))$, together with $\mathcal{M}_{1}(2n)=1\oplus y$, this yields $\mathcal{M}_{1}(4n+y)$. So, if $y=0$, then $\mathcal{M}_{1}(4n)=x=0$, and if $y=1$, then $\mathcal{M}_{1}(4n+1)=1-x=1$, so that $b(4n+1)=y$. Similarly, if $x=1$, then $b(4n+1)=\mathcal{M}_{1}(8n+3)=\mathcal{M}_{1}(4(2n)+3)=1-b(2n)$. With $\mathcal{M}_{1}(2n)=y$, $b(2n)=\mathcal{M}_{1}(4n+1-y)$, we find that if $y=0$, then $1-\mathcal{M}_{1}(4n+1)=x=1$, and if $y=1$, then $1-\mathcal{M}_{1}(4n)=1-x=0$, so that $b(4n+1)=1\oplus y$. Consequently, we obtain that $s(4n+1)=(1\oplus x,\;x\oplus y)$.
Case $r=2$. The first coordinate is $\mathcal{M}_{1}(4n+2)=b(n)=y$. For the second coordinate, $b(4n+2)=\mathcal{M}_{1}(8n+5-y)$. If $y=0$, then $\mathcal{M}_{1}(8n+5)=\mathcal{M}_{1}(4(2n+1)+1)=1-\mathcal{M}_{1}(2n+1)=1\oplus(x\oplus y)=1\oplus x$. If $y=1$, then $\mathcal{M}_{1}(8n+4)=\mathcal{M}_{1}(4(2n+1))=\mathcal{M}_{1}(2n+1)=x\oplus y=1\oplus x$. In both cases, we see that $s(4n+2)=(y,\;1\oplus x)$.
Case $r=3$. The first coordinate is $\mathcal{M}_{1}(4n+3)=1-y=1\oplus y$. For the second coordinate, $b(4n+3)=\mathcal{M}_{1}(8n+6+y)$. If $y=0$, then $\mathcal{M}_{1}(8n+6)=\mathcal{M}_{1}(4(2n+1)+2)=b(2n+1)$. Since $\mathcal{M}_{1}(2n+1)=x$, $b(2n+1)=\mathcal{M}_{1}(4n+3-x)$. So, if $x=0$, then $\mathcal{M}_{1}(4n+3)=1\oplus y=1$; if $x=1$, then $\mathcal{M}_{1}(4n+2)=y=0$; so $b(4n+3)=x\oplus y\oplus 1$. If $y=1$, then $\mathcal{M}_{1}(8n+7)=\mathcal{M}_{1}(4(2n+1)+3)=1-b(2n+1)$. Since $\mathcal{M}_{1}(2n+1)=x\oplus 1$, we find that $b(2n+1)=\mathcal{M}_{1}(4n+2+x)$. So, if $x=0$, then $1-\mathcal{M}_{1}(4n+2)=1-y=0$; if $x=1$, then $1-\mathcal{M}_{1}(4n+3)=y=1$, so that $b(4n+3)=x\oplus y\oplus 1$. Thus $s(4n+3)=(1\oplus y,\;x\oplus y\oplus 1)$. ∎
Theorem 21 (DFAO for $\mathcal{M}_{1}$).
The sequence $\mathcal{M}_{1}$ is $4$-automatic and is computed (MSB-first, base $4$) by the DFAO with states $\{0,1,2,3\}$, initial state 0, output map $\tau(0)=\tau(2)=0$, $\tau(1)=\tau(3)=1$, and transition table
$\delta$ 0 1 2 3
0 0 1 1 2
1 1 2 3 0
2 2 3 0 1
3 3 0 2 3
Proof.
Lemma 20 gives four $\mathbb{F}_{2}$-affine transition rules:
$(x,y)$  $\xrightarrow{\;0\;}(x,\;y),$ $(x,y)$  $\xrightarrow{\;1\;}(1\oplus x,\;x\oplus y),$
$(x,y)$  $\xrightarrow{\;2\;}(y,1\oplus x),$ $(x,y)$  $\xrightarrow{\;3\;}(1\oplus y,\;x\oplus y\oplus 1).$
Encoding the states as $\mathsf{0}=(0,1)$, $\mathsf{1}=(1,1)$, $\mathsf{2}=(0,0)$, $\mathsf{3}=(1,0)$, these yield the stated transition table. The initial state is $s(0)=(\mathcal{M}_{1}(0),\,b(0))=(0,\,\mathcal{M}_{1}(1))=(0,1)$, corresponding to state $\mathsf{0}$. The output reads the first coordinate $x=\mathcal{M}_{1}(n)$. ∎
0/00/01/11/12/02/03/13/10000,30,31,21,2111133332233222211
Figure 2: DFAO for $\mathcal{M}_{1}$ (4 states, base 4, MSB-first). State labels are state/output. The initial state is $\mathsf{0}$.
Remark 22.
The DFAO of $\mathcal{M}_{1}$ has $\mathbb{F}_{2}$-affine transitions, as does that of $\mathcal{M}_{2}$ (Section 4). All three sequences $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ arise from the same mechanism: The balance constraint turns meta-automatic selectors into $\mathbb{F}_{2}$-affine maps on the lifted state space.
Theorem 23.
Neither $\mathcal{M}_{1}$ nor $\mathcal{M}_{2}$ is denestable: In each case, the closure of the $4$-kernel over $\mathbb{Z}$ requires a lifted state of dimension $2$. Moreover, both sequences are $\mathbb{F}_{2}$-linearizable, with
$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n),$
for $\mathcal{M}_{1}$, and
$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n)\oplus\mathcal{M}_{2}(n),\qquad\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1)\oplus\mathcal{M}_{2}(n).$
Proof.
The $\mathbb{F}_{2}$-linearization follows directly from Lemma 10: Since $\mathcal{M}_{1}$ is balanced, we see that $\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$, giving a fixed-argument recurrence.
For non-denestability, suppose the $4$-kernel could be closed using $\mathcal{M}_{1}(n)$ alone, without a second state component. The auxiliary sequence $b(n)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))$ equals $\mathcal{M}_{1}(2n)$ when $\mathcal{M}_{1}(n)=1$ and $\mathcal{M}_{1}(2n+1)$ when $\mathcal{M}_{1}(n)=0$. Since $\mathcal{M}_{1}(2n)\neq\mathcal{M}_{1}(2n+1)$ by balance, no single fixed argument $cn+d$ over $\mathbb{Z}$ reproduces both cases, as $b$ is not of the form $\mathcal{M}_{1}(cn+d)$. Thus the $4$-kernel cannot be closed over $\mathbb{Z}$ without the lifted state, confirming that $\mathcal{M}_{1}$ is not denestable. ∎
Remark 24.
This stands in contrast to Allouche and Shallit [4], who show that the frequency sequence $F(n)$ of the Hofstadter variant $V=Q_{1,4}$ satisfies a recurrence of the form F(2a)=g(F(a2),F(a1),F(a),F(a+1))F(2a)=g(F(a-2),F(a-1),F(a),F(a+1)) with fixed shifts, making $F$ directly denestable in our sense (Definition 12). In our setting, the sequences themselves (not derived quantities) are meta-automatic from the start, and it is the balance constraint that recovers automaticity via $\mathbb{F}_{2}$-linearization rather than denestability.
Corollary 25 (Uniform morphism for $\mathcal{M}_{1}$).
The coding $\tau\colon\mathsf{0},\mathsf{2}\mapsto 0$ and $\mathsf{1},\mathsf{3}\mapsto 1$ applied to the fixed point of the $4$-uniform morphism
$\mathsf{0}\to\mathsf{0112},\quad\mathsf{1}\to\mathsf{1230},\quad\mathsf{2}\to\mathsf{2301},\quad\mathsf{3}\to\mathsf{3023}$
starting from $\mathsf{0}$ produces the sequence $\mathcal{M}_{1}$.
Proof.
For each state $q$, the morphism image is $\delta(q,0)\,\delta(q,1)\,\delta(q,2)\,\delta(q,3)$, read from the transition table of Theorem 21. ∎
4 Meta(2)
Definition 26.
Define $\mathcal{M}_{2}:\mathbb{N}_{0}\to\{0,1\}$ by $\mathcal{M}_{2}(0)=0$, $\mathcal{M}_{2}(1)=1$, and for $n\geq 0$, so that
$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n)),\ \mathcal{M}_{2}(4n+1)=1-\mathcal{M}_{2}(4n)$ (12)
and
$\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-\mathcal{M}_{2}(n)),\ \mathcal{M}_{2}(4n+3)=1-\mathcal{M}_{2}(4n+2).$ (13)
Lemma 27.
The sequence $\mathcal{M}_{2}$ is well-defined and satisfies (4).
Proof.
Well-definedness follows by induction as in Lemma 18, since the selector indices satisfy $2n+\mathcal{M}_{2}(n)\leq 2n+1<4n$ and $2n+1-\mathcal{M}_{2}(n)\leq 2n+1<4n+2$. Balancedness is immediate from the complementary definitions in (12)–(13). ∎
4.1 Lifted state and affine closure
Define the lifted state
$s(n):=\bigl(\mathcal{M}_{2}(n),\ \mathcal{M}_{2}(2n+\mathcal{M}_{2}(n))\bigr)\in\mathbb{F}_{2}^{2}.$ (14)
Lemma 28.
If $s(n)=(x,y)$, then
$\mathcal{M}_{2}(2n)=x\oplus y,\qquad\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1.$
Proof.
If $x=0$, then $y=\mathcal{M}_{2}(2n)$; if $x=1$, then $y=\mathcal{M}_{2}(2n+1)$, and balancedness gives that $\mathcal{M}_{2}(2n)=1-y$. The identities follow. ∎
Theorem 29.
Writing $s(n)=(x,y)$, we have
$s(4n)$ $=(y,\ x\oplus y),$
$s(4n+1)$ $=(y\oplus 1,\ x\oplus y),$
$s(4n+2)$ $=(y\oplus 1,\ x\oplus y\oplus 1),$
$s(4n+3)$ $=(y,\ x\oplus y\oplus 1).$
Consequently $\mathcal{M}_{2}$ is $4$-automatic with at most $4$ states.
Proof.
We use Definition 26, Lemma 28, and balancedness throughout. Write $s(n)=(x,y)$ so that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
Case $r=0$. The first coordinate is given by $\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+x)$. By the recovery lemma, we have that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$. In both cases, for $x=0$ or $x=1$, a direct check gives that $\mathcal{M}_{2}(2n+x)=y$. For the second coordinate $\mathcal{M}_{2}(8n+y)$, we see that if $y=0$, then $\mathcal{M}_{2}(8n)=\mathcal{M}_{2}(4(2n))=\mathcal{M}_{2}(2\cdot 2n+\mathcal{M}_{2}(2n))$, and, since $\mathcal{M}_{2}(2n)=x\oplus y=x$, this equals $\mathcal{M}_{2}(4n+x)$. Simplifying both cases gives $x$. If $y=1$, then $\mathcal{M}_{2}(8n+1)=1-\mathcal{M}_{2}(8n)$, and, by the same analysis as before, we get $x\oplus 1$. In both sub-cases the second coordinate equals $x\oplus y$. Therefore, we obtain
$s(4n)=(y,\;x\oplus y).$
Case $r=1$. The first coordinate is given by $\mathcal{M}_{2}(4n+1)=1-\mathcal{M}_{2}(4n)=1-y=y\oplus 1$.
For the second coordinate of $s(4n+1)$, by definition, this is $\mathcal{M}_{2}(2(4n+1)+\mathcal{M}_{2}(4n+1))=\mathcal{M}_{2}(8n+2+(y\oplus 1))$. When $y=0$, we need $\mathcal{M}_{2}(8n+3)=\mathcal{M}_{2}(4(2n)+3)=1-\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$. Since $\mathcal{M}_{2}(2n)=x$, this equals $1-\mathcal{M}_{2}(4n+1-x)=x\oplus y$ for both values of $x$. When $y=1$, we need $\mathcal{M}_{2}(8n+2)=\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$. Since $\mathcal{M}_{2}(2n)=x\oplus 1$, this equals $\mathcal{M}_{2}(4n+x)=x\oplus y$ for both values of $x$. The second coordinate is therefore $x\oplus y$ in both cases. Therefore
$s(4n+1)=(y\oplus 1,\;x\oplus y).$
Case $r=2$. For the first coordinate $\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-x)$, by XOR linearization (Lemma 10), this equals $\mathcal{M}_{2}(2n+1)\oplus x=(x\oplus y\oplus 1)\oplus x=y\oplus 1$.
For the second coordinate of $s(4n+2)$, we find that $\mathcal{M}_{2}(2(4n+2)+\mathcal{M}_{2}(4n+2))=\mathcal{M}_{2}(8n+4+(y\oplus 1))$. Since $\mathcal{M}_{2}(2n+1)=x\oplus 1$, we have $\mathcal{M}_{2}(8n+4)=\mathcal{M}_{2}(4n+3)=1-(y\oplus 1)$ for both values of $x$. When $y=0$, this gives $\mathcal{M}_{2}(8n+4)=0$, so $\mathcal{M}_{2}(8n+5)=1=x\oplus y\oplus 1$. When $y=1$, since $\mathcal{M}_{2}(2n+1)=x$, we get $\mathcal{M}_{2}(8n+4)=\mathcal{M}_{2}(4n+2+x)=x\oplus y\oplus 1$ by a direct case check. In both cases the second coordinate is $x\oplus y\oplus 1$. It then follows that
$s(4n+2)=(y\oplus 1,\;x\oplus y\oplus 1).$
Case $r=3$. The first coordinate is given by $\mathcal{M}_{2}(4n+3)=1-\mathcal{M}_{2}(4n+2)=1-(y\oplus 1)=y$.
For the second coordinate of $s(4n+3)$, we have that $\mathcal{M}_{2}(2(4n+3)+\mathcal{M}_{2}(4n+3))=\mathcal{M}_{2}(8n+6+y)$. Since $\mathcal{M}_{2}(2n+1)=x\oplus 1$, the identity $\mathcal{M}_{2}(8n+6)=\mathcal{M}_{2}(2(2n+1)+1-\mathcal{M}_{2}(2n+1))$ gives $\mathcal{M}_{2}(8n+6)=x\oplus y\oplus 1$ by a direct two-case check on $x$; this handles $y=0$. For $y=1$, since $\mathcal{M}_{2}(2n+1)=x$, the same formula yields $1-\mathcal{M}_{2}(4n+2+x)=x\oplus y\oplus 1$. The second coordinate is $x\oplus y\oplus 1$ in both cases. Consequently, we have that
$s(4n+3)=(y,\;x\oplus y\oplus 1).$
Each transition has the form $(x,y)\mapsto A\binom{x}{y}+\mathbf{v}(d)$ over $\mathbb{F}_{2}$, confirming the affine structure. ∎
Remark 30.
The linear part $(x,y)\mapsto(y,x\oplus y)$ is given by multiplication by the matrix
$A=\begin{pmatrix}0&1\\ 1&1\end{pmatrix}\in\mathrm{GL}_{2}(\mathbb{F}_{2}),$
which has order $3$ (indeed $A^{3}=I$). The translation vectors are $\mathbf{v}(0)=\binom{0}{0}$, $\mathbf{v}(1)=\binom{1}{0}$, $\mathbf{v}(2)=\binom{1}{1}$, $\mathbf{v}(3)=\binom{0}{1}$.
Theorem 31.
The sequence $\mathcal{M}_{2}$ is computed (MSB-first, base $4$) by the DFAO with states $\{0,1,2,3\}$, initial state 0, output $\tau(0)=\tau(2)=0$, $\tau(1)=\tau(3)=1$, and transition table
$\delta$ 0 1 2 3
0 0 1 3 2
1 2 3 1 0
2 3 2 0 1
3 1 0 2 3
where the encoding is $\mathsf{0}=(0,0)$, $\mathsf{1}=(1,0)$, $\mathsf{2}=(0,1)$, $\mathsf{3}=(1,1)$.
Proof.
This follows directly from the affine transitions in Theorem 29 applied to each encoded state. The initial state is $s(0)=(0,0)$, corresponding to state $\mathsf{0}$. ∎
0/00/01/11/12/02/03/13/10221133110033332233112201122
Figure 3: DFAO for $\mathcal{M}_{2}$ (4 states, base 4, MSB-first). State labels are state/output. The initial state is $\mathsf{0}$.
Corollary 32.
The coding $\tau\colon\mathsf{0},\mathsf{2}\mapsto 0$ and $\mathsf{1},\mathsf{3}\mapsto 1$ applied to the fixed point of the $4$-uniform morphism
$\mathsf{0}\to\mathsf{0132},\quad\mathsf{1}\to\mathsf{2310},\quad\mathsf{2}\to\mathsf{3201},\quad\mathsf{3}\to\mathsf{1023}$
starting from $\mathsf{0}$ produces the sequence $\mathcal{M}_{2}$.
Remark 33.
The four letters of the lifted alphabet correspond to the states $(x,y)\in\mathbb{F}_{2}^{2}$ of the affine automaton (Theorem 29) via
$\mathsf{0}=(0,0),\quad\mathsf{1}=(0,1),\quad\mathsf{2}=(1,0),\quad\mathsf{3}=(1,1).$
With this identification, the morphism $\sigma_{\mathcal{M}_{2}}$ is derived directly from the affine rule $s\mapsto As+\mathbf{v}(d)$, since the image $\sigma_{\mathcal{M}_{2}}(\mathsf{r})$ lists the four successive states visited when the current state is $\mathsf{r}$, for digits $d\in\{0,1,2,3\}$ processed in order. The coding $\tau(\mathsf{r})=x_{\mathsf{r}}\bmod 2$ (first coordinate of the state) then recovers $\mathcal{M}_{2}$. Under the symmetric labeling $a=\mathsf{0}$, $b=\mathsf{1}$, $c=\mathsf{2}$, $d=\mathsf{3}$, the morphism reads $a\to acdb$, $b\to dbac$, $c\to bdca$, $d\to cabd$, with $\tau(a)=\tau(b)=0$ and $\tau(c)=\tau(d)=1$. This alternative notation—where the alphabet labels match the parity of the binary output—is sometimes more convenient for right-special factor analysis.
4.2 An explicit digit formula
Write the base-$4$ expansion $n=\sum_{k\geq 0}d_{k}4^{k}$ with digits $d_{k}\in\{0,1,2,3\}$, and write $d_{k}=2j_{k}+i_{k}$ with $i_{k},j_{k}\in\{0,1\}$.
Theorem 34.
With the notation above, we have that $\mathcal{M}_{2}(n)$ satisfies
$\mathcal{M}_{2}(n)=\bigoplus_{k\equiv 0\ (\mathrm{mod}\ 3)}(i_{k}\oplus j_{k})\ \oplus\ \bigoplus_{k\equiv 1\ (\mathrm{mod}\ 3)}j_{k}\ \oplus\ \bigoplus_{k\equiv 2\ (\mathrm{mod}\ 3)}i_{k}.$
Proof.
The mapping $s\mapsto As+\mathbf{v}(d)$ with $A^{3}=I$ (see Theorem 29 and Remark 30) gives
$s(n)=\sum_{j=0}^{L-1}A^{L-1-j}\,\mathbf{v}(d_{j})\pmod{2}$
after starting from $s(0)=(0,0)$ and processing $L$ base-44 digits MSB-first. Since $A^{3}=I$, the exponent reduces modulo 33. Computing $A^{k}\mathbf{v}(d)$ for $k\in\{0,1,2\}$ and each digit $d$, we obtain
$\begin{array}[]{c|ccc}d&A^{0}\mathbf{v}(d)&A^{1}\mathbf{v}(d)&A^{2}\mathbf{v}(d)\\ \hline\cr 0&(0,0)&(0,0)&(0,0)\\ 1&(1,0)&(0,1)&(1,1)\\ 2&(1,1)&(1,0)&(0,1)\\ 3&(0,1)&(1,1)&(1,0)\end{array}$
Extracting the first coordinate and identifying $i_{k}=d_{k}\bmod 2$ and $j_{k}=\lfloor d_{k}/2\rfloor$ yields the stated formula. ∎
Theorem 35.
Write the binary expansion $n=\sum_{m\geq 0}b_{m}2^{m}$ with $b_{m}\in\{0,1\}$, and define
$q(n):=\sum_{\begin{subarray}{c}m\geq 0\\ m\ (\mathrm{mod}\ 6)\in\{0,1,3,4\}\end{subarray}}b_{m}2^{m},$ (15)
so that $q(n)$ is obtained from $n$ by zeroing the bits in positions $m\equiv 2,5\pmod{6}$. Then
$\mathcal{M}_{2}(n)=\mathbf{t}(q(n))\qquad(n\geq 0).$
Equivalently, if $n=\sum_{k\geq 0}d_{k}4^{k}$ with $d_{k}=2j_{k}+i_{k}$, then $q(n)=\sum_{k\geq 0}d^{\prime}_{k}4^{k}$ where
$d^{\prime}_{k}=\begin{cases}d_{k},&\text{if $k\equiv 0\pmod{3}$;}\\ 2j_{k},&\text{if $k\equiv 1\pmod{3}$;}\\ i_{k},&\text{otherwise.}\end{cases}$
and $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$.
Proof.
In Theorem 34, each base-$4$ digit $d_{k}$ contributes (via $i_{k},j_{k}$) to $\mathcal{M}_{2}(n)$ in a way that depends only on $k\bmod 3$: For $k\equiv 0$, we take $i_{k}\oplus j_{k}$, for $k\equiv 1$ we take only $j_{k}$, and for $k\equiv 2$ we take only $i_{k}$. Since $i_{k}$ and $j_{k}$ are precisely the binary digits $b_{2k}$ and $b_{2k+1}$, this means that $\mathcal{M}_{2}(n)$ is the XOR of all binary digits of $n$ except those in positions $2k$ with $k\equiv 1\pmod{3}$ and $2k+1$ with $k\equiv 2\pmod{3}$, i.e., except $m\equiv 2,5\pmod{6}$. Thus $\mathcal{M}_{2}(n)$ is the parity of the binary digit-sum of $q(n)$ defined in (15), hence $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$. The base-$4$ description of $q(n)$ is the same statement grouped into pairs of bits. ∎
5 Factor complexity
Let $p_{a}(n)$ denote the number of distinct length-$n$ factors in the infinite word $a(0)a(1)a(2)\cdots$.
Lemma 36.
For every infinite binary word $a$ and every $n\geq 1$,
$p_{a}(n+1)-p_{a}(n)=\#\mathrm{RS}_{a}(n),$
where $\mathrm{RS}_{a}(n)$ denotes the set of length-$n$ factors $u$ of $a$ such that both $u0$ and $u1$ are factors of $a$.
Proof.
For each length-$n$ factor $u$, let $E_{R}(u)\subseteq\{0,1\}$ be the set of right extensions such that $u\alpha$ occurs in $a$. Then $p_{a}(n+1)=\sum_{u}|E_{R}(u)|$ and $p_{a}(n)=\sum_{u}1$, so $p_{a}(n+1)-p_{a}(n)=\sum_{u}(|E_{R}(u)|-1)$. In the binary case, $|E_{R}(u)|-1$ equals $1$ if $u$ is right-special and 0 otherwise. ∎
Lemma 37 (Regularity of the right-special count).
If $a$ is $k$-automatic, then the function $n\mapsto\#\mathrm{RS}_{a}(n)$ is $k$-regular in the sense of Allouche–Shallit [3], hence effectively computable from any DFAO for $a$. When $\mathrm{RS}_{a}(n)$ is additionally bounded, it is $k$-automatic.
Proof.
The predicate “$u$ is a length-$n$ factor of aa” and the predicate “$u$ is right-special” are first-order definable over $(\mathbb{N},+,V_{k})$ relative to the DFAO for $a$, so the set of encodings of such pairs is regular and recognized by a finite automaton. Counting right-special factors of a given length reduces to a finite-automaton computation, giving a $k$-regular sequence. For details and background, we again refer to the appropriate texts from Allouche and Shallit [3] and from Shallit [27]. If the values are bounded, $k$-regularity reduces to $k$-automaticity. ∎
Several claims in this section can be mechanically certified using the Walnut theorem prover [27], which decides first-order properties of automatic sequences in Presburger arithmetic. For clarity, and in the hope of developing a deeper understanding of the behavior of meta-automatic sequences, we present “non-Walnut” proofs wherever appropriate.
5.1 Quarto
We first record a key property of the Thue–Morse factor language.
Lemma 38.
The language $\mathrm{Fac}(\mathbf{t})$ is closed under bitwise complementation, i.e., if $w\in\mathrm{Fac}(\mathbf{t})$, then $\bar{w}\in\mathrm{Fac}(\mathbf{t})$.
Proof.
The Thue–Morse morphism $\mu:0\mapsto 01,\;1\mapsto 10$ satisfies $\overline{\mu(w)}=\mu(\bar{w})$ for every binary word $w$. Since $\mathbf{t}=\mu^{\omega}(0)$ and both 0 and $1=\bar{0}$ appear in $\mathbf{t}$, every factor $w$ of $\mu^{\omega}(0)$ yields $\bar{w}=\overline{\mu^{\omega}(0)}|_{\text{same position}}=\mu^{\omega}(1)|_{\text{same position}}$, which is a factor of $\mu^{\omega}(1)=1\,\mathbf{t}[1\,\infty)$, hence a factor of $\mathbf{t}$. ∎
Proposition 39.
Let $J_{n}$ denote the set of length-$n$ factors of $\mathcal{Q}$ that are not factors of $\mathbf{t}$. Then
$p_{\mathcal{Q}}(n)=p_{\mathbf{t}}(n)+|J_{n}|.$
Proof.
By Proposition 16, on each dyadic block $2^{k},2^{k+1})$ the sequence $\mathcal{Q}$ coincides with $\mathbf{t}$ or $1-\mathbf{t}$ according to the parity of $k$. A factor of $\mathcal{Q}$ of length $n$ either lies entirely within one block or straddles a dyadic boundary.
Intra-block factors. If the factor lies in an even block it is a factor of $\mathbf{t}$. If it lies in an odd block it is a complement of a factor of $\mathbf{t}$, which by Lemma 38 is again a factor of $\mathbf{t}$. So all intra-block factors belong to $\mathrm{Fac}(\mathbf{t})$.
Junction factors. A factor straddling boundary $2^{k}$ has the form $u\cdot\bar{v}$ (or $\bar{u}\cdot v$, by a symmetric argument), where $u\cdot v\in\mathrm{Fac}(\mathbf{t})$. Such a word need not lie in $\mathrm{Fac}(\mathbf{t})$: If $u\cdot\bar{v}\in\mathrm{Fac}(\mathbf{t})$, then $\bar{u}\cdot v\in\mathrm{Fac}(\mathbf{t})$ by Lemma 38, but this need not hold in general. We define $J_{n}$ as the set of those length-$n$ junction factors not already in $\mathrm{Fac}(\mathbf{t})$.
Since every factor of $\mathbf{t}$ appears in $\mathcal{Q}$ (via even blocks, which contain $\mathbf{t}$ itself), the factor language of $\mathcal{Q}$ is exactly $\mathrm{Fac}(\mathbf{t})\cup J_{n}$, giving $p_{\mathcal{Q}}(n)=p_{\mathbf{t}}(n)+|J_{n}|$. ∎
Theorem 40.
For all $k\geq 4$, we have that
$p_{\mathcal{Q}}(2^{k})=13\cdot 2^{k-2}-2.$
Equivalently, $p_{\mathcal{Q}}(2^{k})=p_{\mathbf{t}}(2^{k})+2^{k-2}$, so $|J_{2^{k}}|=2^{k-2}$.
Proof.
The value $p_{\mathbf{t}}(2^{k})=3\cdot 2^{k}-2$ is known [6, 23]. By Proposition 39, $p_{\mathcal{Q}}(2^{k})=p_{\mathbf{t}}(2^{k})+|J_{2^{k}}|$. The right-special count $\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n)$ takes values in $\{2,3,4\}$, verified computationally for $2\leq n\leq 3000$ (using 30 00030\,000 terms of $\mathcal{Q}$); a complete proof for all $n$ can be obtained through a Walnut certificate, obtainable from the DFAO for $\mathcal{Q}$. It follows from the recurrence
$p_{\mathcal{Q}}(2^{k+2})=4\,p_{\mathcal{Q}}(2^{k})+6\qquad(k\geq 4),$
with $p_{\mathcal{Q}}(16)=50$, which has the unique solution $p_{\mathcal{Q}}(2^{k})=13\cdot 2^{k-2}-2$ and which we verify computationally for $4\leq k\leq 7$. ∎
Remark 41.
The junction set $J_{2^{k}}$ has size $2^{k-2}=\Theta(2^{k})$: Junction factors account for exactly $\tfrac{1}{4}$ of all length-$2^{k}$ factors of $\mathcal{Q}$ beyond those of $\mathbf{t}$.
Since $\mathrm{RS}_{\mathcal{Q}}(n)\in\{2,3,4\}$ for all $n\geq 1$, the values of $p_{\mathcal{Q}}$ between dyadic points are completely determined. The support of each slope takes a precise regular form.
Proposition 42.
For $m\geq 4$, set $k=\lfloor\log_{2}n\rfloor$ and write $n=2^{k}+i$ with $0\leq i<2^{k}$. Then
$p_{\mathcal{Q}}(n)=\begin{cases}13\cdot 2^{k-2}-2,&\text{if $i=0$;}\\ 4n-3\cdot 2^{k-2}-4,&\text{if $1\leq i\leq 2^{k-1}$;}\\ 3n+3\cdot 2^{k-2}-3,&\text{if $2^{k-1}<i\leq 3\cdot 2^{k-2}$;}\\ 2n+5\cdot 2^{k-1}-2,&\text{if $3\cdot 2^{k-2}<i<2^{k}$.}\end{cases}$
In other words, on each block $2^{k},2^{k+1})$, the right-special count $\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n)$ equals $4$ for $n\in(2^{k},\,3\cdot 2^{k-1}]$, then $3$ for $n\in(3\cdot 2^{k-1},\,7\cdot 2^{k-2}]$, then $2$ for $n\in(7\cdot 2^{k-2},\,2^{k+1})$. The formulas agree at the breakpoints and are continuous across dyadic boundaries. In particular, Theorem 40 is the restriction to $i=0$.
Proof.
A complete proof for all $n$ can be obtained via Walnut, leaving the details to the reader. ∎
As a consequence of Proposition 42, we have that we have $n+1\leq p_{\mathcal{Q}}(n)=O(n)$ for all $n\geq 1$.
5.2 Meta(1) and Meta(2)
Proposition 43.
We have $p_{\mathcal{M}_{1}}(n)=\Theta(n)$ and $p_{\mathcal{M}_{2}}(n)=\Theta(n)$.
Proof.
The incidence matrices of the morphisms in Corollaries 25 and 32 are
$M_{\mathcal{M}_{1}}=\begin{pmatrix}1&2&1&0\\ 1&1&1&1\\ 1&1&1&1\\ 1&0&1&2\end{pmatrix},\qquad M_{\mathcal{M}_{2}}=\begin{pmatrix}1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\end{pmatrix}$.
One checks $M_{\mathcal{M}_{1}}^{2}>0$ and $M_{\mathcal{M}_{2}}>0$ directly, so both morphisms are primitive. From the work of Pansiot [24], the fixed point of a primitive $k$-uniform morphism over an alphabet $\mathcal{A}$ has complexity $p(n)=\Theta(n)$. Each of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is a coding $\tau(\sigma^{\omega}(a_{0}))$, where $\tau\colon\mathcal{A}\to\{0,1\}$ is non-eventually-constant. Since every letter appears with positive frequency and $\tau$ takes both values, distinct factors of $\sigma^{\omega}(a_{0})$ not identified by $\tau$ remain distinct in $\tau(\sigma^{\omega}(a_{0}))$, giving $p_{\tau(\sigma^{\omega})}(n)\geq p_{\sigma^{\omega}}(n)/|\mathcal{A}|^{2}=\Theta(n)$ [3, Proposition 10.9.2]. ∎
The exact complexity at dyadic lengths, and the complete piecewise formula, are given in the theorems below.
Example 44.
We obtain the initial values
$p_{\mathcal{Q}}(1\ldots 15)=(2,4,6,10,12,16,20,23,26,30,34,38,42,45,48),$
and
$p_{\mathcal{M}_{2}}(1\ldots 15)=(2,4,6,10,12,14,16,18,20,24,28,32,36,40,44).$
Proposition 45.
For all $k\geq 3$,
$p_{\mathcal{M}_{2}}(2^{k})=\begin{cases}\dfrac{5}{2}\cdot 2^{k}-2,&\text{if $3\mid k$;}\\ 3\cdot 2^{k}-2,&\text{otherwise.}\end{cases}$
In particular, the dyadic slope alternates between $\frac{5}{2}$ and $3$ according to $k\bmod 3$. The three cases modulo $3$ are tied together by the recurrence
$p_{\mathcal{M}_{2}}(2^{k+3})=8\,p_{\mathcal{M}_{2}}(2^{k})+14\qquad(k\geq 3).$
Proof.
The formula is certifiable by the Walnut theorem prover, leaving the details to the reader. ∎
Since $\mathrm{RS}_{\mathcal{M}_{2}}(n)\in\{2,4\}$, the values of $p_{\mathcal{M}_{2}}$ between dyadic points are determined by the support of $\mathrm{RS}_{\mathcal{M}_{2}}=4$. This support has a precise structure that extends Theorem 45 to a complete evaluation.
Proposition 46.
For $n\geq 1$, set $k=\lfloor\log_{2}n\rfloor$ and $r=k\bmod 3$.
Case $r=2$: On the entire block $n\in[2^{k},2^{k+1}]$,
$p_{\mathcal{M}_{2}}(n)=2n+(2^{k}-2).$
Case $r=0$: The block $n\in[2^{k},2^{k+1}]$ splits into three linear pieces,
$p_{\mathcal{M}_{2}}(n)=\begin{cases}2n+(2^{k-1}-2),&\text{if $2^{k}\leq n\leq 2^{k}+1$;}\\ 4n-(3\cdot 2^{k-1}+4),&\text{if $2^{k}+1\leq n\leq 7\cdot 2^{k-2}+1$;}\\ 2n+(2^{k+1}-2),&\text{if $7\cdot 2^{k-2}+1\leq n\leq 2^{k+1}$.}\end{cases}$
Case $r=1$: The block $n\in[2^{k},2^{k+1}]$ splits into three linear pieces,
$p_{\mathcal{M}_{2}}(n)=\begin{cases}2n+(2^{k}-2),&\text{if $2^{k}\leq n\leq 2^{k}+1$;}\\ 4n-(2^{k}+4),&\text{if $2^{k}+1\leq n\leq 3\cdot 2^{k-1}+1$;}\\ 2n+(2^{k+1}-2),&\text{if $3\cdot 2^{k-1}+1\leq n\leq 2^{k+1}$.}\end{cases}$
Proof.
Again, a proof certificate can be obtained via Walnut, leaving the details to the reader. ∎
6 Concluding remarks
A parallel direction replaces the dyadic balancedness given by $a(2n)+a(2n+1)=1$ by the ternary constraint $a(3n)+a(3n+1)+a(3n+2)=2$, while still restricting to binary values in $\{0,1\}$. This forces exactly two $1$s and one 0 in each base-$3$ block, giving density $2/3$ on prefixes of length $3^{m}$. Meta-automatic selectors of the form a(3n+ra(n))a(3n+r-a(n)) arise naturally, and the $\mathbb{F}_{2}$-linearization mechanism of Theorem 23 extends to this setting: while such selectors are not denestable, the ternary balance constraint may again convert them into fixed-argument $\mathbb{F}_{2}$-affine maps, yielding $3$-automatic sequences. A small number of initial configurations appear to produce binary $3$-automatic sequences with small DFAOs; exploring this family and its complexity functions is left for a separate treatment.
7 Appendix
Minimal DFAO sizes for the meta-automatic sequences considered in this paper are listed in Table 1.
Table 1: Minimal DFAO size for each specialization. “Sel.” counts the number of meta-automatic selectors among $F$ and $G$. Boldface marks the three sequences studied in this paper.
$F(n)$ $G(n)$ DFAO Sel. Notes
$a(n)$ $a(n)$ 22 0 Thue–Morse
$a(n)$ $a(2n)$ $\mathbf{3}$ $\mathbf{0}$ Quarto $\mathcal{Q}$
$a(n)$ $a(2n+1)$ 22 0 Thue–Morse
$a(n)$ $a(2n+a(n))$ 44 11
$a(n)$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{1}$ Meta(1) $\mathcal{M}_{1}$
$a(2n)$ $a(n)$ 44 0
$a(2n)$ $a(2n)$ 0 ultimately periodic
$a(2n)$ $a(2n+1)$ 22 0 Thue–Morse
$a(2n)$ $a(2n+a(n))$ 33 11
$a(2n)$ a(2n+1a(n))a(2n+1-a(n)) 33 11
$a(2n+1)$ $a(n)$ 55 0
$a(2n+1)$ $a(2n)$ 33 0 complement of $\mathcal{Q}$
$a(2n+1)$ $a(2n+1)$ 55 0
$a(2n+1)$ $a(2n+a(n))$ 55 11
$a(2n+1)$ a(2n+1a(n))a(2n+1-a(n)) 55 11
$a(2n+a(n))$ $a(n)$ 44 11
$a(2n+a(n))$ $a(2n)$ 44 11
$a(2n+a(n))$ $a(2n+1)$ 44 11
$a(2n+a(n))$ $a(2n+a(n))$ 44 22
$a(2n+a(n))$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{2}$ Meta(2) $\mathcal{M}_{2}$
a(2n+1a(n))a(2n+1-a(n)) $a(n)$ 44 11
a(2n+1a(n))a(2n+1-a(n)) $a(2n)$ 44 11
a(2n+1a(n))a(2n+1-a(n)) $a(2n+1)$ 44 11
a(2n+1a(n))a(2n+1-a(n)) $a(2n+a(n))$ 44 22 complement of $\mathcal{M}_{2}$
a(2n+1a(n))a(2n+1-a(n)) a(2n+1a(n))a(2n+1-a(n)) 44 22
8 Acknowledgments
Claude Opus 4.6 (Anthropic) was used as a computational assistant, in particular for exploring the factor-complexity functions of $\mathcal{Q}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{1}$ of Section 5 and guessing the formulas, and systematically investigating variants of the recursions as in Section 7, which allowed us to select $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ as three sequences of particular interest for our purposes. The first author thanks Jean-Paul Allouche, Karl Dilcher, and Jeffrey Shallit for useful feedback related to this paper.

### v2

@@
 This inspires us to introduce what may be seen as a natural combination of the recurrences for meta-Fibonacci and automatic sequences, by introducing the concept of a meta-automatic sequence.
-We show how it is possible to construct a meta-automatic sequence that is not denestable in terms of not being reducible, in a specific sense formalized in this paper, to certain digit-based recurrences for automatic sequences.
-This motivates our study of the non-denestable, meta-automatic sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ introduced in this paper.
-For each of these integer sequences, we prove explicit DFAO evaluations, together with $4$-uniform morphisms, and we also consider the factor complexities of these sequences.
+We exhibit two binary meta-automatic sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ whose defining recurrences do not satisfy the Allouche–Shallit automaticity criterion directly, and this is formalized in our paper.
+For each of these integer sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, we prove explicit DFAO evaluations, together with 4-uniform morphisms, and we also consider the factor complexities of these sequences.
 1 Introduction
-Informally, a sequence $(a(n))_{n\geq 0}$ is said to satisfy a nested or meta-Fibonacci recurrence if $a(n)$, for all sufficiently large $n$, can be expressed in terms of $a$ evaluated at an expression involving a value given as a term in $(a(n))_{n\geq 0}$.
-The classical and prototypical example is Hofstadter’s $Q$-sequence [19, p.
-137], which is indexed in the On-Line Encyclopedia of Integer Sequences [28] as A005185, and which satisfies $Q(1)=Q(2)=1$ and
+A nested recurrence is a recurrence in which a term $a(n)$ is defined using values of $a$ at indices that themselves depend on earlier values of the sequence.
+The classical example is Hofstadter’s $Q$-sequence [20, p.
+137], indexed in the On-Line Encyclopedia of Integer Sequences [30] as A005185, defined by $Q(1)=Q(2)=1$ and
 Q(n)=Q(nQ(n1))+Q(nQ(n2))Q(n)=Q(n-Q(n-1))+Q(n-Q(n-2)) (1)
 for $n>2$.
-In contrast to nested recurrences, digit-based recurrences (which may also be referred to as divide-and-conquer recurrences) define $a\big(k^{t}n+j\big)$ (for $j\in\{0,1,\ldots,k^{t}-1\}$) as a function of previously computed values of the form $a\big(k^{t^{\prime}}n^{\prime}+j^{\prime}\big)$ for $t^{\prime}<t$ (and for $j^{\prime}\in\{0,1,\ldots,k^{t^{\prime}}-1\}$).
-These digit-based recurrences are used to define automatic sequences [3, 9], which are of great importance in many areas of number theory, computer science, and combinatorics.
-The depth of the study of meta-Fibonacci sequences is evidenced by the many open problems related to Hofstadter’s $Q$-sequence and generalizations.
-Notably, to this day, it is not even known whether or not $Q(n)$ is defined for all integers $n>0$.
-This together with the roles played by automatic sequences within many disciplines in mathematics and computer science lead us to consider a new and interdisciplinary subject that may be thought of as “combining” the recursions for meta-Fibonacci and automatic sequences.
-As it turns out, relationships between meta-Fibonacci and automatic sequences have previously been considered.
-Notably, Allouche and Shallit [4], in 2012, showed that the frequency sequence of a variant of Hofstadter’s $Q$-sequence satisfies a digit-based recurrence, so that this variant sequence is $2$-automatic.
-In this context, they proved an automaticity criterion [4, Theorem 2.2] reviewed in Section 1.2 below.
-In this paper, we introduce the concept of a meta-automatic sequence, which, informally, can be thought of as extending both the Allouche–Shallit criterion and the recursions for meta-Fibonacci sequences as in (1).
+Recurrences as in (1) are also known as meta-Fibonacci recurrences.
+Other examples of integer sequences satisfying nested recurrences include the sequence A244477 studied by Golomb [18].
+Many problems concerning Hofstadter’s $Q$-sequence and its generalizations remain open to this day.
+Notably, it is not known whether or not $Q(n)$ is even defined for all $n>0$.
+In contrast to nested recurrences, digit-based recurrences (also called divide-and-conquer recurrences) define $a(k^{t}n+j)$ for $j\in\{0,1,\ldots,k^{t}-1\}$ as a function of previously computed values $a(k^{t^{\prime}}n^{\prime}+j^{\prime})$ with $t^{\prime}<t$ and $j^{\prime}\in\{0,1,\ldots,k^{t^{\prime}}-1\}$.
+Digit-based recurrences can be used to define automatic sequences [3, 9], which appear in many areas of number theory, computer science, and combinatorics.
+Allouche and Shallit [4] showed that the frequency sequence of a variant of Hofstadter’s $Q$-sequence satisfies a digit-based recurrence, and, moreover, is $2$-automatic.
+They also proved an automaticity criterion [4, Theorem 2.2] reviewed in Section 1.2 below.
+This inspires us to introduce the concept of a meta-automatic sequence, which may, informally, be thought of as “combining” the concepts of meta-Fibonacci and automatic sequences.
+This is formalized below.
 For example, in contrast to a rule such as $a(4n+2)=a(2n)$ used to define an automatic sequence, we consider rules such as
 a(4n+2)=a(2n+1a(n)),a(4n+2)=a(2n+1-a(n)), (2)
-where the argument of the right-hand side depends on the value $a(n)$.
-Recurrences as in (2) do not, in general, fall under the Allouche–Shallit automaticity criterion [4, Theorem 2.2], since the argument on the right-hand side of (2) is not necessarily equivalent to a polynomial in $n$ of the desired form.
-The idea of combining the recursions associated with automatic and meta-Fibonacci sequences is motivated by many past references concerning meta-Fibonacci sequences, in addition to Allouche and Shallit’s work on the automaticity of the frequency sequence for Hofstadter’s $Q$-sequence [4].
-In this direction, we highlight notable references on the Hofstadter $Q$-sequence [16, 17, 18, 25], and on variants and generalizations of the Hofstadter $Q$-sequence [1, 2, 5, 12, 14, 30] and on meta-Fibonacci sequences [7, 8, 10, 11, 13, 15, 20, 21, 22, 26, 29].
+in which the argument of $a$ on the right-hand side depends on the value $a(n)$.
+A recurrence of the form suggested in (2) would not, in general, satisfy the Allouche–Shallit automaticity criterion, since the argument on the right is not of the required polynomial form.
+The idea of combining automatic and meta-Fibonacci recurrences is motivated by past work on meta-Fibonacci sequences, in addition to the above referenced work by Allouche and Shallit [4].
+We point to references on the Hofstadter $Q$-sequence [16, 17, 19, 27], on variants and generalizations of the Hofstadter $Q$-sequence [1, 2, 5, 12, 14, 32], and on meta-Fibonacci sequences [7, 8, 10, 11, 13, 15, 21, 22, 23, 28, 31].
 1.1 Outline
-A formal definition of the term meta-automatic sequence is given in Section 2 below.
-As we demonstrate, it is possible for the recurrence for a meta-automatic sequence to, in a specific sense, degenerate in such a way so that it reduces to a digit-based recurrence.
-To formalize what should be meant by such reductions, we introduce the concept of denestability.
-As it turns out, there exist meta-automatic sequences that are not denestable, and we introduce the first construction of a sequence of this form.
-In this direction, Theorem 23 shows that the integer sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ that we introduce in this paper are not denestable.
-This leads to the following question motivating the topic of our paper: Under what conditions do meta-automatic recurrences produce automatic sequences?
-The binary sequences $(a(n))_{n\geq 0}$ under consideration in this paper are mainly balanced in the sense that $a(2n)+a(2n+1)=1$.
-This condition provides a defining property of the Thue–Morse sequence, and so the meta-automatic sequences introduced in this paper may be seen as “meta-Thue–Morse” sequences.
-We prove that in base $2$, meta-automatic sequences that are binary and balanced are ultimately periodic (Lemma 14).
-However, as we later clarify, by passing to a base-$4$ digit template that preserves a dyadic balance, this can be used to construct balanced, binary, and meta-automatic sequences that are not ultimately periodic.
-In this direction, a key idea is given by how the balance constraint linearizes each meta-automatic selector via an $\mathbb{F}_{2}$-affine operation (Lemma 10).
-This motivates our study of the following meta-automatic sequences:
-  • Meta(1) $\mathcal{M}_{1}$ (A392736): a mixed template in which one of the two branch rules uses a meta-automatic selector. The XOR linearization yields a $4$-state DFAO (Section 3).
-  • Meta(2) $\mathcal{M}_{2}$ (A391614): a “pure” two-selector template in which both branch rules are meta-automatic.
-The linearized recurrence has an affine structure over $\mathbb{F}_{2}^{2}$ governed by the matrix $\bigl(\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\bigr)$, and the sequence admits the factorization $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, where $q$ is an explicit bit-masking operator and $\mathbf{t}$ denotes the Thue–Morse sequence (Theorem 35).
-Remarkably, despite $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ sharing the same automaton size of $4$ states, the two sequences differ sharply in structural complexity.
-Notably, we find that $\mathcal{M}_{2}$ admits the evaluation $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, while $\mathcal{M}_{1}$ requires a lifted size-two kernel.
-Despite this shared automaton size, their factor-complexity functions, studied in Section 5, differ markedly.
-For each of the above sequences, we give an MSB-first, base-$4$ DFAO, an explicit $4$-uniform morphism, and factor-complexity results.
-The broader theme is that the interplay between balancedness and $\mathbb{F}_{2}$-linearity can bring meta-automatic, digit-based recurrences into the realm of automatic sequences, even when criteria for automaticity, such as the above referenced Allouche–Shallit criterion [4], do not apply.
+A formal definition of a meta-automatic sequence is given in Section 2.
+A meta-automatic recurrence may, after simplification, reduce to a digit-based recurrence, in which case the Allouche–Shallit criterion applies directly.
+We formalize, in this paper, how meta-automatic sequences are not necessarily reducible in this way.
+A binary sequence $a:\mathbb{N}_{0}\to\{0,1\}$ is said to be balanced if
+$a(2n)+a(2n+1)=1\qquad(n\geq 0).$ (3)
+The binary sequences under consideration in this paper are, for the most part, balanced.
+In a related way, much of this paper makes use of the additive operation underlying the field $\mathbb{F}_{2}=\{0,1\}$, letting $\oplus$ denote this operation, which may be referred to as XOR, with $0\oplus 0=1\oplus 1=0$ and $0\oplus 1=1\oplus 0=1$.
+For balanced binary sequences, Lemma 5 gives the identity
+$a(2n+1-a(n))=a(2n+1)\oplus a(n).$
+We prove that balanced, binary, meta-automatic sequences are, for the cases covered in Lemma 6, ultimately periodic.
+To obtain non-periodic examples, we therefore work with base-$4$ recurrences.
+We consider:
+  • The integer sequence $\mathcal{M}_{1}$ (A392736), defined via a base-$4$ recurrence in which one of the rules in the recurrence is nested in the sense outlined above.
+After applying Lemma 5 below, the pair $(\mathcal{M}_{1}(n),b(n))$, for $b(n)$ defined in (8) below, yields a 44-state DFAO (Section 3).
+  • The integer sequence $\mathcal{M}_{2}$ (A391614), defined via a base-$4$ recurrence in which both of the rules in the recurrence are nested in the sense outlined above.
+After applying Lemma 5, the pair $\bigl(\mathcal{M}_{2}(n),\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n))\bigr)$ obeys an affine rule over $\mathbb{F}_{2}^{2}$ with linear part $\bigl(\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\bigr)$.
+The sequence satisfies $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, where $q$ is an explicit bit-masking operator and $\mathbf{t}$ is the Thue–Morse sequence (Theorem 26).
+Although $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ both have $4$-state DFAOs, we have that $\mathcal{M}_{2}$ admits the evaluation $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$, whereas $\mathcal{M}_{1}$ is described by means of the pair $(\mathcal{M}_{1}(n),b(n))$.
+Their factor-complexity functions, studied in Section 5, also differ.
+For each of the sequences $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, we give a base-$4$ DFAO (MSB-first) and an explicit $4$-uniform morphism, and we also investigate the factor complexity functions for $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in Section 5.
 1.2 Preliminaries
-For background on automatic sequences, we refer to the appropriate texts from Allouche and Shallit [3] and from Shallit [27].
-We briefly recall some definitions that are useful or neeed, for our purposes.
-In this direction, the concept of a $k$-kernel [3, §6.6] is required in our work.
+For background on automatic sequences, we refer to the texts of Allouche and Shallit [3] and of Shallit [29].
+We proceed with the following definition, which may be used, as below, to define automatic sequences.
 Definition 1.
@@
 $K_{k}(U)=\left\{\left(U_{k^{i}n+j}\right)_{n\geq 0}:\text{$i\geq 0$ and $0\leq j<k^{i}$}\right\}.$
-For $k\geq 2$, a $k$-automatic sequence may then be defined as a sequence $U=(U(n))_{n\geq 0}$ such that $K_{k}(U)$ is finite.
-Equivalently, this sequence is $k$-automatic if there exists a DFAO that, on input the base-$k$ representation of $n$, outputs $U(n)$, again referring to Allouche and Shallit’s text for background.
-Definition 2.
-Fix a base $k\geq 2$.
-By a digit-based recurrence we mean a system of recurrences that defines $a(kn+r)$ for digits $r\in\{0,1,\ldots,k-1\}$ in terms of values of $a(m)$ at strictly smaller indices $m<kn+r$ (typically among $m=n,2n,\dots$).
-Let $\mathbf{t}:\mathbb{N}_{0}\to\{0,1\}$ denote the Thue–Morse sequence, defined by $\mathbf{t}(0)=0$ and
+For $k\geq 2$, a $k$-automatic sequence is a sequence $U=(U(n))_{n\geq 0}$ such that $K_{k}(U)$ is finite.
+Equivalently, the sequence $U$ is $k$-automatic if there exists a DFAO such that, upon inputting the base-$k$ representation of $n$, this DFAO outputs $U(n)$.
+See Allouche and Shallit’s text [3] for details.
+A prototypical instance of an automatic sequence is the Thue–Morse sequence $\mathbf{t}:\mathbb{N}_{0}\to\{0,1\}$ defined so that $\mathbf{t}(0)=0$ and
 $\mathbf{t}(2n)=\mathbf{t}(n),\qquad\mathbf{t}(2n+1)=1-\mathbf{t}(n)\qquad(n\geq 0).$
-Equivalently, $\mathbf{t}(n)=s_{2}(n)\bmod 2$, where $s_{2}(n)$ is the sum of the binary digits of $n$.
-Since much of our work is closely related to an automaticity criterion due to Allouche and Shallit [4], it is necessary, for our purposes, to review this criterion, as below.
+Equivalently, $\mathbf{t}(n)=s_{2}(n)\bmod 2$, where $s_{2}(n)$ is the sum of the binary digits of $n$, noting that the balanced condition in (3) holds.
+We introduce, in this paper, a variant of an automaticity criterion due to Allouche and Shallit [4], which we now recall.
 Let $U=(U(n))_{n\geq 0}$ denote a sequence, and let $\alpha$ denote an integer.
@@
 Moreover, for integers $q\geq 2$ and $t\geq 1$, we write $s(q,t)=\frac{1-q^{t+1}}{1-q}$ in place of the number of pairs $(i,j)$ of integers $i$ and $j$ such that $0\leq i\leq t$ and $0\leq j\leq q^{i}-1$.
-We then order expressions of the form $U_{q,i,j}$ by writing $U_{1}=U$, $U_{2}$, $\ldots$, $U_{s(q,t)}$.
+Fix an ordering of these pairs $(i,j)$, with $(0,0)$ appearing first, and write the corresponding sequences as $U_{1}=U,\,U_{2},\ldots,\,U_{s(q,t)}$.
+The choice of ordering plays no role in what follows.
 We also let $m(q,t)=q^{t+1}-1$.
 Theorem 2.
-(Allouche &\& Shallit, 2012) Let the terms of $(U(n))_{n\geq 0}$ be in a finite set $\mathcal{A}$, and let $q\geq 2$ denote an integer.
+(Allouche and Shallit, 2012) Let the terms of $(U(n))_{n\geq 0}$ be in a finite set $\mathcal{A}$, and let $q\geq 2$ denote an integer.
 Then $U$ is $q$-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ together with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions from $\mathcal{A}^{a+b+s(q,t)}$ to $\mathcal{A}$ such that
@@
 2 Meta-automatic sequences
-Let $q>0$ and $i_{0}>0$ and $j_{0}\geq 0$ denote integers.
-For some $v\geq 0$, and for indices $k\in\{1,2,\ldots,v\}$, let $\mathcal{S}_{k}$ and $i_{k}>0$ and $j_{k}\geq 0$ be integers, subject to the condition that $i_{k}<i_{0}$ for all indices $k$.
-By analogy with (3), and with the assumption that the codomain of $U$ is contained in $\mathbb{N}_{0}$, we write
-$U_{q,i_{0},j_{0},\mathcal{S}_{1},i_{1},j_{1},\ldots,\mathcal{S}_{v},i_{v},j_{v}}=U\Big(q^{i_{0}}n_{0}+j_{0}+\mathcal{S}_{1}U(q^{i_{1}}n_{1}+j_{1})+\cdots+\mathcal{S}_{v}U(q^{i_{v}}n_{v}+j_{v})\Big).$
+Let $q>0$, $i_{0}>0$, and $j_{0}\geq 0$ be integers.
+For some $v\geq 0$, and for indices $k\in\{1,2,\ldots,v\}$, let $c_{k}$, $i_{k}>0$, and $j_{k}\geq 0$ be integers, with $i_{k}<i_{0}$ for all $k$.
+By analogy with (4), and assuming that the codomain of $U$ is contained in $\mathbb{N}_{0}$, we write
+$U_{q,i_{0},j_{0},c_{1},i_{1},j_{1},\ldots,c_{v},i_{v},j_{v}}(n)=U\Big(q^{i_{0}}n+j_{0}+c_{1}U(q^{i_{1}}n+j_{1})+\cdots+c_{v}U(q^{i_{v}}n+j_{v})\Big).$
+The following definition may be thought of as formalizing the intuition behind recurrences as in (2).
 Definition 3.
-Let the terms of $(U(n))_{n\geq 0}$ be in a finite subset $\mathcal{A}$ of $\mathbb{N}_{0}$.
-Let $q\geq 2$ be an integer.
-Then $U$ is $q$-meta-automatic if there are nonnegative integers $t$, $a$, $b$, and $n_{0}$ along with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions to $\mathcal{A}$ (with domains given below) such that, for all $n\geq n_{0}$ and for all $j\in\{0,1,\ldots,m(q,t)\}$, the value $U\big(q^{t+1}n+j\big)$ is equal to $f_{j}$ evaluated at a tuple of the form
-$\Big(U^{-a}(n),\ldots,U^{-1}(n),U^{0}(n),U^{1}(n),\ldots,U^{b}(n),\\ U_{q,i_{0}^{(1)},j_{0}^{(1)},\mathcal{S}_{1}^{(1)},i_{1}^{(1)},j_{1}^{(1)},\ldots,\mathcal{S}_{v_{1}}^{(1)},i_{v_{1}}^{(1)},j_{v_{1}}^{(1)}},\ldots,U_{q,i_{0}^{(w)},j_{0}^{(w)},\mathcal{S}_{1}^{(w)},i_{1}^{(w)},j_{1}^{(w)},\ldots,\mathcal{S}_{v_{w}}^{(w)},i_{v_{w}}^{(w)},j_{v_{w}}^{(w)}}\Big)$
-for some $w\geq 0$ and for integers $i_{k_{1}}^{(k_{2})}\leq t$ and $j_{k_{1}}^{(k_{2})}$ and $\mathcal{S}_{k_{1}}^{(k_{2})}$, and where the domain of $f_{j}$ is $\mathcal{A}^{a+b+w+1}$.
+Suppose that each term of $(U(n))_{n\geq 0}$ is in a finite subset $\mathcal{A}$ of $\mathbb{N}_{0}$, and let $q\geq 2$ be an integer.
+The sequence $U$ is $q$-meta-automatic if there exist nonnegative integers $t$, $a$, $b$, $n_{0}$ together with a family $\{f_{j}:j=0,1,\ldots,m(q,t)\}$ of functions to $\mathcal{A}$, with domains as below, such that, for all $n\geq n_{0}$ and all $j\in\{0,1,\ldots,m(q,t)\}$, the value $U(q^{t+1}n+j)$ equals $f_{j}$ evaluated at the tuple
+$\Big(U^{-a}(n),\,\ldots,\,U^{-1}(n),\,U^{0}(n),\,U^{1}(n),\,\ldots,\,U^{b}(n),\\ U_{q,i_{0}^{(1)},j_{0}^{(1)},c_{1}^{(1)},i_{1}^{(1)},j_{1}^{(1)},\ldots,c_{v_{1}}^{(1)},i_{v_{1}}^{(1)},j_{v_{1}}^{(1)}}(n),\,\ldots,\\ U_{q,i_{0}^{(w)},j_{0}^{(w)},c_{1}^{(w)},i_{1}^{(w)},j_{1}^{(w)},\ldots,c_{v_{w}}^{(w)},i_{v_{w}}^{(w)},j_{v_{w}}^{(w)}}(n)\Big)$
+for some $w\geq 0$ and integers $i_{k_{1}}^{(k_{2})}\leq t$, $j_{k_{1}}^{(k_{2})}$, and $c_{k_{1}}^{(k_{2})}$, where the domain of $f_{j}$ is $\mathcal{A}^{a+b+w+1}$.
 Example 4.
-Let the integer sequence $(\mathcal{T}(n))_{n\geq 0}$ be defined so that $\mathcal{T}(0)=0$ and so that
+Define $(\mathcal{T}(n))_{n\geq 0}$ by $\mathcal{T}(0)=0$ and
 $\mathcal{T}(2n)$ $=\mathcal{T}(n-\mathcal{T}(n-1))\ \text{for $n\geq 1$, and}$
 $\mathcal{T}(2n+1)$ $=1-\mathcal{T}(n)\ \text{for $n\geq 0$}.$
-According to Definition 4, the sequence $(\mathcal{T}(n))_{n\geq 0}$ is 22-meta-automatic.
-It can be shown that this sequence agrees, for $n\geq 1$, with the OEIS sequence A039982, which is defined in a different way and without the use of meta-recursions.
-Definition 6.
-For a $q$-meta-automatic sequence $(U(n))_{n\geq 0}$, adopting notation from Definition 4, a selector refers to an expression given by $f_{j}$ evaluated at the tuple displayed in Definition 4, i.e., with such an expression being equal to $U(q^{t+1}n+j)$.
-Example 7.
-For the $2$-meta-automatic sequence in Example 5, the expression $\mathcal{T}(n-\mathcal{T}(n-1))$ is a selector.
-Binary sequences of the following form are a main object of study in our work.
-Definition 8.
-A binary sequence $a:\mathbb{N}_{0}\to\{0,1\}$ is said to be balanced if
-$a(2n)+a(2n+1)=1\qquad(n\geq 0).$ (4)
-Example 9.
-The Thue–Morse sequence may be seen as a prototypical instance of a balanced (binary) sequence.
-As in the following result, we let $\oplus$ denote the additive operation in the field $\mathbb{F}_{2}$.
+Then $(\mathcal{T}(n))_{n\geq 0}$ is $2$-meta-automatic in the sense of Definition 3. For $n\geq 1$, this sequence agrees with the OEIS sequence A039982, which was defined without nested recurrences.
 Lemma 5.
@@
 Proof.
-This follows in a direct way from (4). ∎
-Remark 11.
-The selector a(2n+1a(n))a(2n+1-a(n)) and the XOR expression $a(2n+1)\oplus a(n)$ compute the same function of the triple $(a(n),a(2n),a(2n+1))$.
-Both read the same two independent bits $(a(n),a(2n))$, since balancedness gives a(2n+1)=1a(2n)a(2n+1)=1-a(2n).
-The balance constraint does not reduce the number of value dependencies, but it ensures that the dependency is $\mathbb{F}_{2}$-affine: The selector, viewed as a Boolean function, is addition in $\mathbb{F}_{2}$.
-2.1 Denestability
-We introduce two notions that distinguish the sequences studied in this paper from those covered by the automaticity criterion of Allouche and Shallit [4, Theorem 2.2].
-Definition 12.
-A meta-automatic recurrence for a sequence $(U(n))_{n\geq 0}$ is denestable if every selector associated with an expression of the form $U\big(q^{t+1}n+j\big)$ can, for sufficiently large $n\geq n_{0}$, be written as an expression of the form given in the right-hand side of the identity in Theorem 3, without increasing the size of kernel.
-Example 13.
-The integer sequence $(\mathcal{T}(n))_{n\geq 0}$ defined in Example 5 is denestable, since the given recursion is such that $\mathcal{T}(2n)=1$ for all $n\geq 1$, providing, together with the relation $\mathcal{T}(2n+1)=1-\mathcal{T}(n)$, a digit-based recurrence for $\mathcal{T}$.
-A weaker notion, relative to Definition 12, is that of $\mathbb{F}_{2}$-linearizability, which, informally, allows the balance constraint to convert meta-automatic selectors into $\mathbb{F}_{2}$-affine maps at fixed arguments.
-This is later clarified.
-In base $2$, the most direct way to introduce a meta-automatic argument while keeping the binary balancedness (4) is, informally, to try to define the even subsequence using a “one-level” selector at the $n$-scale.
-For instance, set
+The first identity follows from (3) by case analysis on $a(n)\in\{0,1\}$, and similarly for the second. ∎
+2.1 Base-4 recurrences
+We now record a result that motivates the use of base-$4$ recurrences in the sequel. We let
 $a(2n)=a(n+a(n))\qquad\text{or}\qquad a(2n)=a(n+1-a(n)),$
-and then set a(2n+1)=1a(2n)a(2n+1)=1-a(2n) to preserve (4). Under balance, however, both schemes collapse to ultimate periodicity, and this is formalized below.
+and we let a(2n+1)=1a(2n)a(2n+1)=1-a(2n). In both cases, we obtain ultimate periodicity, as we now show.
 Lemma 6.
@@
   1. (a)
-    If $a(2n)=a(n+a(n))$ for all $n\geq 0$, then $a(0)=0$ and
-    $a(2n)=0,\qquad a(2n+1)=1\qquad(n\geq 2),$
-    so $a$ is ultimately periodic of period $2$.
+    If $a(2n)=a(n+a(n))$ for all $n\geq 0$, then $a(0)=0$ and $a$ is ultimately $2$-periodic, with
+    $a(2n)=0,\qquad a(2n+1)=1\qquad(n\geq 2).$
   2. (b)
-    If a(2n)=a(n+1a(n))a(2n)=a(n+1-a(n)) for all $n\geq 0$, then $a(0)=1$ and
-    $a(2n)=1,\qquad a(2n+1)=0\qquad(n\geq 2),$
-    so $a$ is ultimately periodic of period $2$.
+    If a(2n)=a(n+1a(n))a(2n)=a(n+1-a(n)) for all $n\geq 0$, then $a(0)=1$ and $a$ is ultimately $2$-periodic, with
+    $a(2n)=1,\qquad a(2n+1)=0\qquad(n\geq 2).$
 Proof.
-(a) Taking $n=0$ gives $a(0)=a(a(0))$. If $a(0)=1$, then balancedness forces $a(1)=0$, and hence $a(0)=a(1)=0$, a contradiction. Thus $a(0)=0$ and $a(1)=1$.
-Write $b_{m}=a(2m)$, so that balancedness gives $a(2m+1)=1-b_{m}$. Using $a(2n)=a(n+a(n))$ with $n=2m$, we obtain
-$b_{2m}=a(4m)=a(2m+a(2m))=a(2m+b_{m})=\begin{cases}a(2m)=b_{m},&b_{m}=0;\\[2.0pt] a(2m+1)=1-b_{m}=0,&b_{m}=1.\end{cases}$
+(a) Taking $n=0$ gives $a(0)=a(a(0))$. If $a(0)=1$, then balancedness gives us that $a(1)=0$, and hence $a(0)=a(1)=0$, a contradiction. Thus $a(0)=0$ and $a(1)=1$.
+Write $b_{m}=a(2m)$, so that balancedness gives that $a(2m+1)=1-b_{m}$. Using $a(2n)=a(n+a(n))$ with $n=2m$, we obtain
+$b_{2m}=a(4m)=a(2m+a(2m))=a(2m+b_{m})=\begin{cases}a(2m)=b_{m},&\text{if }b_{m}=0;\\ a(2m+1)=1-b_{m}=0,&\text{if }b_{m}=1.\end{cases}$
 So, we have that $b_{2m}=0$ for all $m\geq 0$.
-Next, using $a(2n)=a(n+a(n))$ with $n=2m+1$ and $a(2m+1)=1-b_{m}$, we get
-$b_{2m+1}=a(4m+2)=a(2m+1+a(2m+1))=a(2m+2-b_{m})=\\ \begin{cases}a(2m+2)=b_{m+1},&b_{m}=0;\\[2.0pt] a(2m+1)=0,&b_{m}=1.\end{cases}$
+Now, using $a(2n)=a(n+a(n))$ with $n=2m+1$ and $a(2m+1)=1-b_{m}$, we get
+$b_{2m+1}=a(4m+2)=a(2m+1+a(2m+1))=a(2m+2-b_{m})=\\ \begin{cases}a(2m+2)=b_{m+1},&\text{if }b_{m}=0;\\ a(2m+1)=0,&\text{if }b_{m}=1.\end{cases}$
 We claim that $b_{m}=0$ for all $m\geq 2$.
@@
 (b) This follows from (a) by complementing: if $a$ satisfies (b) then $\bar{a}(n)=1-a(n)$ satisfies (a), and vice versa. ∎
-Lemma 14 shows that meta-automatic behavior requires passing to base 44: the dyadic pairing (4) is preserved, but there is now room for selectors one scale below.
-2.2 A non-nested baseline case
-Fix $a(0)=0$, $a(1)=1$, and define
+Lemma 6 leads us to consider base-44 recurrences such that the balance identity in (3) is preserved.
+2.2 A non-nested case
+Fix $a(0)=0$ and $a(1)=1$, and define
 $a(4n)=F(n),\quad a(4n+1)=1-F(n),a(4n+2)=G(n),\quad a(4n+3)=1-G(n),$ (5)
 so that $a(2n)+a(2n+1)=1$ holds automatically.
-The three main sequences investigated in this paper arise by specializing (5).
-To begin with, we consider an integer sequence $(\mathcal{Q}(n))_{n\geq 0}$ that does not have a nested structure in terms of its recurrence but that is to be used extensively in our subsequent work.
+The sequences studied in this paper arise as specializations of (5).
+We first consider a sequence $(\mathcal{Q}(n))_{n\geq 0}$ whose recurrence is not nested, but which is used in the sequel.
 Definition 7.
 The Thue–Morse Quarto sequence $\mathcal{Q}:\mathbb{N}_{0}\to\{0,1\}$ (A298952, where we use the convention $\mathcal{Q}(0)=0$ rather than $a(0)=1$ as in the OEIS entry) is defined by $\mathcal{Q}(0)=0$, $\mathcal{Q}(1)=1$, and
-$\mathcal{Q}(4n)$ $=\mathcal{Q}(n),$ $\mathcal{Q}(4n+1)$ $=1-\mathcal{Q}(n),$ (6)
-$\mathcal{Q}(4n+2)$ $=\mathcal{Q}(2n),$ $\mathcal{Q}(4n+3)$ $=1-\mathcal{Q}(2n).$ (7)
-Omitting details for brevity, we have that there is a $3$-state DFAO for $(\mathcal{Q}(n))_{n\geq 0}$, in such a way so that $\mathcal{Q}$ is $4$-automatic and computed by the DFAO with states $\{A,B,C\}$, initial state $A$, and output $\tau(A)=\tau(B)=0$, $\tau(C)=1$.
-This DFAO is shown in Figure 1.
-The DFAO for $\mathcal{Q}$ corresponds to the $4$-uniform morphism
+$\mathcal{Q}(4n)$ $=\mathcal{Q}(n),$ $\mathcal{Q}(4n+1)$ $=1-\mathcal{Q}(n),$
+$\mathcal{Q}(4n+2)$ $=\mathcal{Q}(2n),$ $\mathcal{Q}(4n+3)$ $=1-\mathcal{Q}(2n).$
+The sequence $\mathcal{Q}$ is $4$-automatic.
+It is computed by the $3$-state DFAO with states in $\{A,B,C\}$, initial state $A$, and output $\tau(A)=\tau(B)=0$, $\tau(C)=1$, shown in Figure 1.
+This DFAO corresponds to the $4$-uniform morphism
 $A\mapsto ACBC,\quad B\mapsto BCCB,\quad C\mapsto CBBC,$
-with coding $\tau(A)=\tau(B)=0$, $\tau(C)=1$.
+along with a coding such that $\tau(A)=\tau(B)=0$ and $\tau(C)=1$.
 A/0A/0B/0B/0C/1C/100,30,30,30,3221,31,31,21,21,21,2
@@
 Proposition 8.
-Define $d:\mathbb{N}_{0}\to\{0,1\}$ by $d(0)=0$ and $d(n)=\lfloor\log_{2}n\rfloor\bmod 2$ for $n\geq 1$. Then
+Define $d:\mathbb{N}_{0}\to\{0,1\}$ so that $d(0)=0$ and $d(n)=\lfloor\log_{2}n\rfloor\bmod 2$ for $n\geq 1$. Then
 $\mathcal{Q}(n)=\mathbf{t}(n)\oplus d(n)\qquad(n\geq 0).$
@@
 Case $r=3$: $\mathcal{Q}(4n+3)=1\oplus\mathcal{Q}(2n)=\mathbf{t}(n)\oplus d(n)$, and $\mathbf{t}(4n+3)\oplus d(4n+3)=\mathbf{t}(n)\oplus d(n)$. ∎
-3 Meta(1)
-Definition 17 (Meta(1)).
-Define $\mathcal{M}_{1}:\mathbb{N}_{0}\to\{0,1\}$ by $\mathcal{M}_{1}(0)=0$, $\mathcal{M}_{1}(1)=1$, and for $n\geq 0$, so that
-$\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n),\ \mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(n),$ (8)
-and
-$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),\ \mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (9)
+3 The sequence $\mathcal{M}_{1}$
+Definition 9.
+Define $\mathcal{M}_{1}(0)=0$ and $\mathcal{M}_{1}(1)=1$. For $n\geq 1$, set
+$\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n),\qquad\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(n).$ (6)
+For $n\geq 0$, set
+$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),\qquad\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (7)
 Lemma 10.
@@
 Well-definedness follows by induction on $m$.
-For $m=4n+r$, the right-hand sides in (8) reduce to indices <m<m; for (9), the selector index is $2n+1-\mathcal{M}_{1}(n)\in\{2n,2n+1\}<4n+2\leq m$.
-Balancedness follows from (5), with $\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(4n)$ and $\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(4n+2)$.
+For $m=4n+r$, the right-hand sides in (6) reduce to indices <m<m; for (7), the nested argument $2n+1-\mathcal{M}_{1}(n)\in\{2n,2n+1\}<4n+2\leq m$.
+Balancedness follows from (5), since $\mathcal{M}_{1}(4n+1)=1-\mathcal{M}_{1}(4n)$ and $\mathcal{M}_{1}(4n+3)=1-\mathcal{M}_{1}(4n+2)$.
 ∎
 Remark 11.
-Because $\mathcal{M}_{1}$ is balanced, Lemma 10 yields
+Since $\mathcal{M}_{1}$ is balanced, Lemma 5 gives
 $\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n),$
 so (7) is equivalent to the non-nested relation $\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$.
-The selector form is the natural “meta” description and is the form that generalizes.
 3.1 A 4-state DFAO
-Define the auxiliary function
-$b(n):=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)).$ (10)
-Consider the lifted state $s(n):=(\mathcal{M}_{1}(n),b(n))\in\{0,1\}^{2}$.
+Define
+$b(n):=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n)),$ (8)
+and consider the pair $s(n):=(\mathcal{M}_{1}(n),b(n))$.
+Informally, the following result gives us that the pair $s(4n+r)$ depends only on $s(n)$, for $r\in\{0,1,2,3\}$ and for $n\geq 1$, disregarding the trivial $n=0$ case whereby $\mathcal{M}_{1}(0)=0$.
 Lemma 12.
-For each $r\in\{0,1,2,3\}$, the state $s(4n+r)$ depends only on $s(n)$.
+Let $x=\mathcal{M}_{1}(n)$ and $y=b(n)$, and let $n\geq 1$. Then
+$s(4n)$ $=(x,y),$ (9)
+$s(4n+1)$ $=(1\oplus x,\;x\oplus y),$ (10)
+$s(4n+2)$ $=(y,\;1\oplus x),\text{and}$ (11)
+$s(4n+3)$ $=(1\oplus y,\;x\oplus y\oplus 1).$ (12)
 Proof.
-Write $s(n)=(x,y)$ where $x=\mathcal{M}_{1}(n)$ and $y=b(n)$.
-By the balancedness given by Lemma 18 and the XOR linearization (Lemma 10), we have $y=b(n)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)=(1-\mathcal{M}_{1}(2n))\oplus x$, so that
-$\mathcal{M}_{1}(2n)=1\oplus x\oplus y,\qquad\mathcal{M}_{1}(2n+1)=x\oplus y.$ (11)
+By Lemma 10 and Lemma 5, we have $y=b(n)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)=(1-\mathcal{M}_{1}(2n))\oplus x$, so that
+$\mathcal{M}_{1}(2n)=1\oplus x\oplus y,\qquad\mathcal{M}_{1}(2n+1)=x\oplus y.$
 Case $r=0$.
 The first coordinate is $\mathcal{M}_{1}(4n)=\mathcal{M}_{1}(n)=x$.
-For the second coordinate, $b(4n)=\mathcal{M}_{1}(8n+1-\mathcal{M}_{1}(4n))=\mathcal{M}_{1}(8n+1-x)$.
-If $x=0$, then $b(4n)=\mathcal{M}_{1}(8n+1)=\mathcal{M}_{1}(4(2n)+1)=1-\mathcal{M}_{1}(2n)=1-(1\oplus y)=y$.
+The latter coordinate satisfies $b(4n)=\mathcal{M}_{1}(8n+1-\mathcal{M}_{1}(4n))=\mathcal{M}_{1}(8n+1-x)$.
+If $x=0$, then $b(4n)=\mathcal{M}_{1}(8n+1)=\mathcal{M}_{1}(4(2n)+1)=1-\mathcal{M}_{1}(2n)=1-(1\oplus x\oplus y)=1-(1\oplus y)=y$.
 If $x=1$, then $b(4n)=\mathcal{M}_{1}(8n)=\mathcal{M}_{1}(4(2n))=\mathcal{M}_{1}(2n)=1\oplus x\oplus y=y$.
-In both cases, we find that s(4n)=(x,y)s(4n)=(x,\,y).
+In both cases, we find that (9) holds.
 Case $r=1$.
 The first coordinate is $\mathcal{M}_{1}(4n+1)=1-x=1\oplus x$.
-For the second coordinate, we have that $b(4n+1)=\mathcal{M}_{1}(8n+2+x)$.
+The latter coordinate satisfies $b(4n+1)=\mathcal{M}_{1}(8n+2+x)$.
 If $x=0$, then $b(4n+1)=\mathcal{M}_{1}(8n+2)=\mathcal{M}_{1}(4(2n)+2)=b(2n)$.
-Using the property that $b(2n)=\mathcal{M}_{1}(4n+1-\mathcal{M}_{1}(2n))$, together with $\mathcal{M}_{1}(2n)=1\oplus y$, this yields $\mathcal{M}_{1}(4n+y)$.
+Using the property that $b(2n)=\mathcal{M}_{1}(4n+1-\mathcal{M}_{1}(2n))$, together with $\mathcal{M}_{1}(2n)=1\oplus y$, this yields $b(2n)=\mathcal{M}_{1}(4n+y)$.
 So, if $y=0$, then $\mathcal{M}_{1}(4n)=x=0$, and if $y=1$, then $\mathcal{M}_{1}(4n+1)=1-x=1$, so that $b(4n+1)=y$.
@@
 With $\mathcal{M}_{1}(2n)=y$, $b(2n)=\mathcal{M}_{1}(4n+1-y)$, we find that if $y=0$, then $1-\mathcal{M}_{1}(4n+1)=x=1$, and if $y=1$, then $1-\mathcal{M}_{1}(4n)=1-x=0$, so that $b(4n+1)=1\oplus y$.
-Consequently, we obtain that $s(4n+1)=(1\oplus x,\;x\oplus y)$.
+Consequently, we obtain that (10) holds.
 Case $r=2$.
 The first coordinate is $\mathcal{M}_{1}(4n+2)=b(n)=y$.
-For the second coordinate, $b(4n+2)=\mathcal{M}_{1}(8n+5-y)$.
+The latter coordinate satisfies $b(4n+2)=\mathcal{M}_{1}(8n+5-y)$.
 If $y=0$, then $\mathcal{M}_{1}(8n+5)=\mathcal{M}_{1}(4(2n+1)+1)=1-\mathcal{M}_{1}(2n+1)=1\oplus(x\oplus y)=1\oplus x$.
 If $y=1$, then $\mathcal{M}_{1}(8n+4)=\mathcal{M}_{1}(4(2n+1))=\mathcal{M}_{1}(2n+1)=x\oplus y=1\oplus x$.
-In both cases, we see that $s(4n+2)=(y,\;1\oplus x)$.
+In both cases, we see that (11) holds.
 Case $r=3$.
 The first coordinate is $\mathcal{M}_{1}(4n+3)=1-y=1\oplus y$.
-For the second coordinate, $b(4n+3)=\mathcal{M}_{1}(8n+6+y)$.
+The latter coordinate satisfies $b(4n+3)=\mathcal{M}_{1}(8n+6+y)$.
 If $y=0$, then $\mathcal{M}_{1}(8n+6)=\mathcal{M}_{1}(4(2n+1)+2)=b(2n+1)$.
@@
 So, if $x=0$, then $1-\mathcal{M}_{1}(4n+2)=1-y=0$; if $x=1$, then $1-\mathcal{M}_{1}(4n+3)=y=1$, so that $b(4n+3)=x\oplus y\oplus 1$.
-Thus $s(4n+3)=(1\oplus y,\;x\oplus y\oplus 1)$.
+This gives us the desired relation in (12).
 ∎
-Theorem 21 (DFAO for $\mathcal{M}_{1}$).
+Theorem 13.
 The sequence $\mathcal{M}_{1}$ is $4$-automatic and is computed (MSB-first, base $4$) by the DFAO with states $\{0,1,2,3\}$, initial state 0, output map $\tau(0)=\tau(2)=0$, $\tau(1)=\tau(3)=1$, and transition table
@@
 Remark 14.
-The DFAO of $\mathcal{M}_{1}$ has $\mathbb{F}_{2}$-affine transitions, as does that of $\mathcal{M}_{2}$ (Section 4).
-All three sequences $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ arise from the same mechanism: The balance constraint turns meta-automatic selectors into $\mathbb{F}_{2}$-affine maps on the lifted state space.
-Theorem 23.
-Neither $\mathcal{M}_{1}$ nor $\mathcal{M}_{2}$ is denestable: In each case, the closure of the $4$-kernel over $\mathbb{Z}$ requires a lifted state of dimension $2$.
-Moreover, both sequences are $\mathbb{F}_{2}$-linearizable, with
-$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n),$
-for $\mathcal{M}_{1}$, and
-$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n)\oplus\mathcal{M}_{2}(n),\qquad\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1)\oplus\mathcal{M}_{2}(n).$
-Proof.
-The $\mathbb{F}_{2}$-linearization follows directly from Lemma 10: Since $\mathcal{M}_{1}$ is balanced, we see that $\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n)$, giving a fixed-argument recurrence.
-For non-denestability, suppose the $4$-kernel could be closed using $\mathcal{M}_{1}(n)$ alone, without a second state component.
-The auxiliary sequence $b(n)=\mathcal{M}_{1}(2n+1-\mathcal{M}_{1}(n))$ equals $\mathcal{M}_{1}(2n)$ when $\mathcal{M}_{1}(n)=1$ and $\mathcal{M}_{1}(2n+1)$ when $\mathcal{M}_{1}(n)=0$.
-Since $\mathcal{M}_{1}(2n)\neq\mathcal{M}_{1}(2n+1)$ by balance, no single fixed argument $cn+d$ over $\mathbb{Z}$ reproduces both cases, as $b$ is not of the form $\mathcal{M}_{1}(cn+d)$.
-Thus the $4$-kernel cannot be closed over $\mathbb{Z}$ without the lifted state, confirming that $\mathcal{M}_{1}$ is not denestable.
-∎
+The four transition rules of the DFAO of $\mathcal{M}_{1}$ are $\mathbb{F}_{2}$-affine in $(x,y)$; the same is true for $\mathcal{M}_{2}$ (Section 4).
+For each of the sequences $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$, the affine form is a consequence of Lemma 5.
+The defining recurrence of $\mathcal{M}_{1}$ is not directly an instance of Theorem 2, since the argument $2n+1-\mathcal{M}_{1}(n)$ in the nested recurrence in (7) depends on the value $\mathcal{M}_{1}(n)$.
+The $b$-sequence defined in (8) removes this dependence.
+Furthermore, Lemma 5 gives us that
+$\mathcal{M}_{1}(4n+2)=\mathcal{M}_{1}(2n+1)\oplus\mathcal{M}_{1}(n).$
+The XOR identity is a direct consequence of Lemma 5, since $\mathcal{M}_{1}$ is balanced.
 Remark 15.
-This stands in contrast to Allouche and Shallit [4], who show that the frequency sequence $F(n)$ of the Hofstadter variant $V=Q_{1,4}$ satisfies a recurrence of the form F(2a)=g(F(a2),F(a1),F(a),F(a+1))F(2a)=g(F(a-2),F(a-1),F(a),F(a+1)) with fixed shifts, making $F$ directly denestable in our sense (Definition 12).
-In our setting, the sequences themselves (not derived quantities) are meta-automatic from the start, and it is the balance constraint that recovers automaticity via $\mathbb{F}_{2}$-linearization rather than denestability.
-Corollary 25 (Uniform morphism for $\mathcal{M}_{1}$).
+By contrast, Allouche and Shallit [4] show that the frequency sequence $F(n)$ of the Hofstadter variant $V=Q_{1,4}$ satisfies a recurrence of the form F(2a)=g(F(a2),F(a1),F(a),F(a+1))F(2a)=g(F(a-2),F(a-1),F(a),F(a+1)) with fixed shifts.
+This satisfies the required conditions in Theorem 2.
+In our setting, the sequences we consider themselves, as opposed to associated frequency sequences, are meta-automatic.
+In this direction, Lemma 5 provides a key tool, for our purposes.
+Corollary 16.
 The coding $\tau\colon\mathsf{0},\mathsf{2}\mapsto 0$ and $\mathsf{1},\mathsf{3}\mapsto 1$ applied to the fixed point of the $4$-uniform morphism
@@
 For each state $q$, the morphism image is $\delta(q,0)\,\delta(q,1)\,\delta(q,2)\,\delta(q,3)$, read from the transition table of Theorem 13. ∎
-4 Meta(2)
+4 The sequence $\mathcal{M}_{2}$
 Definition 17.
-Define $\mathcal{M}_{2}:\mathbb{N}_{0}\to\{0,1\}$ by $\mathcal{M}_{2}(0)=0$, $\mathcal{M}_{2}(1)=1$, and for $n\geq 0$, so that
-$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n)),\ \mathcal{M}_{2}(4n+1)=1-\mathcal{M}_{2}(4n)$ (12)
-and
-$\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-\mathcal{M}_{2}(n)),\ \mathcal{M}_{2}(4n+3)=1-\mathcal{M}_{2}(4n+2).$ (13)
-Lemma 27.
+Define $\mathcal{M}_{2}(0)=0$ and $\mathcal{M}_{2}(1)=1$. For $n\geq 1$, set
+$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n)),\qquad\mathcal{M}_{2}(4n+1)=1-\mathcal{M}_{2}(4n).$ (13)
+For $n\geq 0$, set
+$\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-\mathcal{M}_{2}(n)),\qquad\mathcal{M}_{2}(4n+3)=1-\mathcal{M}_{2}(4n+2).$ (14)
+Theorem 18.
 The sequence $\mathcal{M}_{2}$ is well-defined and satisfies (3).
 Proof.
-Well-definedness follows by induction as in Lemma 18, since the selector indices satisfy $2n+\mathcal{M}_{2}(n)\leq 2n+1<4n$ and $2n+1-\mathcal{M}_{2}(n)\leq 2n+1<4n+2$.
+Well-definedness follows by induction as in Lemma 10.
+For the recurrence rules associated with the arguments $4n$ and $4n+1$, we have $n\geq 1$, so $2n+\mathcal{M}_{2}(n)\leq 2n+1<4n$.
+For the arguments $4n+2$ and $4n+3$, the nested argument satisfies $2n+1-\mathcal{M}_{2}(n)\leq 2n+1<4n+2$.
 Balancedness is immediate from the complementary definitions in (13)–(14).
 ∎
-4.1 Lifted state and affine closure
-Define the lifted state
+Define the pair
 $s(n):=\bigl(\mathcal{M}_{2}(n),\,\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n))\bigr)\in\mathbb{F}_{2}^{2}.$ (15)
@@
 Proof.
-We use Definition 26, Lemma 28, and balancedness throughout. Write $s(n)=(x,y)$ so that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
+The case $n=0$ is immediate from the initial values $\mathcal{M}_{2}(0)=0$ and $\mathcal{M}_{2}(1)=1$.
+We may therefore assume $n\geq 1$ whenever a rule for $4n$ or $4n+1$ is used.
+Write $s(n)=(x,y)$ so that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
 Case $r=0$.
 The first coordinate is given by $\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+x)$.
-By the recovery lemma, we have that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
+By Lemma 19, we have that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
 In both cases, for $x=0$ or $x=1$, a direct check gives that $\mathcal{M}_{2}(2n+x)=y$.
@@
 For the second coordinate of $s(4n+1)$, by definition, this is $\mathcal{M}_{2}(2(4n+1)+\mathcal{M}_{2}(4n+1))=\mathcal{M}_{2}(8n+2+(y\oplus 1))$.
-When $y=0$, we need $\mathcal{M}_{2}(8n+3)=\mathcal{M}_{2}(4(2n)+3)=1-\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$.
+When $y=0$, we have $\mathcal{M}_{2}(8n+3)=\mathcal{M}_{2}(4(2n)+3)=1-\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$.
 Since $\mathcal{M}_{2}(2n)=x$, this equals $1-\mathcal{M}_{2}(4n+1-x)=x\oplus y$ for both values of $x$.
-When $y=1$, we need $\mathcal{M}_{2}(8n+2)=\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$.
+When $y=1$, we have $\mathcal{M}_{2}(8n+2)=\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$.
 Since $\mathcal{M}_{2}(2n)=x\oplus 1$, this equals $\mathcal{M}_{2}(4n+x)=x\oplus y$ for both values of $x$.
@@
 $s(4n+1)=(y\oplus 1,\;x\oplus y).$
-Case $r=2$.
-For the first coordinate $\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-x)$, by XOR linearization (Lemma 10), this equals $\mathcal{M}_{2}(2n+1)\oplus x=(x\oplus y\oplus 1)\oplus x=y\oplus 1$.
+Case $r=2$. For the first coordinate $\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-x)$, Lemma 5 gives $\mathcal{M}_{2}(2n+1)\oplus x=(x\oplus y\oplus 1)\oplus x=y\oplus 1$.
 For the second coordinate of $s(4n+2)$, we find that $\mathcal{M}_{2}(2(4n+2)+\mathcal{M}_{2}(4n+2))=\mathcal{M}_{2}(8n+4+(y\oplus 1))$.
@@
 Remark 24.
-The four letters of the lifted alphabet correspond to the states $(x,y)\in\mathbb{F}_{2}^{2}$ of the affine automaton (Theorem 29) via
-$\mathsf{0}=(0,0),\quad\mathsf{1}=(0,1),\quad\mathsf{2}=(1,0),\quad\mathsf{3}=(1,1).$
-With this identification, the morphism $\sigma_{\mathcal{M}_{2}}$ is derived directly from the affine rule $s\mapsto As+\mathbf{v}(d)$, since the image $\sigma_{\mathcal{M}_{2}}(\mathsf{r})$ lists the four successive states visited when the current state is $\mathsf{r}$, for digits $d\in\{0,1,2,3\}$ processed in order.
-The coding $\tau(\mathsf{r})=x_{\mathsf{r}}\bmod 2$ (first coordinate of the state) then recovers $\mathcal{M}_{2}$.
-Under the symmetric labeling $a=\mathsf{0}$, $b=\mathsf{1}$, $c=\mathsf{2}$, $d=\mathsf{3}$, the morphism reads $a\to acdb$, $b\to dbac$, $c\to bdca$, $d\to cabd$, with $\tau(a)=\tau(b)=0$ and $\tau(c)=\tau(d)=1$.
-This alternative notation—where the alphabet labels match the parity of the binary output—is sometimes more convenient for right-special factor analysis.
+The four letters of the alphabet correspond to the states $(x,y)\in\mathbb{F}_{2}^{2}$ of the affine automaton (Theorem 20) via
+$\mathsf{0}=(0,0),\quad\mathsf{1}=(1,0),\quad\mathsf{2}=(0,1),\quad\mathsf{3}=(1,1).$
+With this identification, the image $\sigma_{\mathcal{M}_{2}}(\mathsf{r})$ lists the four successive states visited when the current state is $\mathsf{r}$ and the digits $d\in\{0,1,2,3\}$ are read in order, and the coding $\tau(\mathsf{r})$ is the first coordinate of the state.
+Under the relabeling $a=\mathsf{0}$, $b=\mathsf{2}$, $c=\mathsf{1}$, $d=\mathsf{3}$, the morphism reads $a\to acdb$, $b\to dbac$, $c\to bdca$, $d\to cabd$, with $\tau(a)=\tau(b)=0$ and $\tau(c)=\tau(d)=1$.
+This labeling is used in the right-special factor analysis below.
+The defining recurrence of $\mathcal{M}_{2}$ is not directly an instance of Theorem 2, since the argument $2n+\mathcal{M}_{2}(n)$ in the nested recurrence in (13) and the argument $2n+1-\mathcal{M}_{2}(n)$ in (14) depend on the value $\mathcal{M}_{2}(n)$.
+This together with the sequence $\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n))$ used in the pair (15) yields the system of recurrences in Theorem 20.
+Furthermore, Lemma 5 gives us that
+$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n)\oplus\mathcal{M}_{2}(n),\qquad\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1)\oplus\mathcal{M}_{2}(n).$
+The XOR identities are direct consequences of Lemma 5, since $\mathcal{M}_{2}$ is balanced.
 4.1 An explicit digit formula
@@
 Theorem 25.
-With the notation above, we have that $\mathcal{M}_{2}(n)$ satisfies
+We have that
 $\mathcal{M}_{2}(n)=\bigoplus_{k\equiv 0\ (\mathrm{mod}\ 3)}(i_{k}\oplus j_{k})\ \oplus\ \bigoplus_{k\equiv 1\ (\mathrm{mod}\ 3)}j_{k}\ \oplus\ \bigoplus_{k\equiv 2\ (\mathrm{mod}\ 3)}i_{k}.$
 Proof.
-The mapping $s\mapsto As+\mathbf{v}(d)$ with $A^{3}=I$ (see Theorem 29 and Remark 30) gives
-$s(n)=\sum_{j=0}^{L-1}A^{L-1-j}\,\mathbf{v}(d_{j})\pmod{2}$
-after starting from $s(0)=(0,0)$ and processing $L$ base-44 digits MSB-first.
-Since $A^{3}=I$, the exponent reduces modulo 33.
-Computing $A^{k}\mathbf{v}(d)$ for $k\in\{0,1,2\}$ and each digit $d$, we obtain
+Let $d_{L-1}\cdots d_{1}d_{0}$ be the base-$4$ expansion of $n$.
+Since the automaton reads digits MSB-first and the transition has the form $s\mapsto As+\mathbf{v}(d)$, starting from $s(0)=(0,0)$ gives
+$s(n)=\sum_{k=0}^{L-1}A^{k}\,\mathbf{v}(d_{k})\pmod{2}$.
+Since $A^{3}=I$, the exponent depends only on $k\bmod 3$. Computing $A^{k}\mathbf{v}(d)$ for $k\in\{0,1,2\}$ and each digit $d$, we obtain
 $\begin{array}[]{c|ccc}d&A^{0}\mathbf{v}(d)&A^{1}\mathbf{v}(d)&A^{2}\mathbf{v}(d)\\ \hline\cr 0&(0,0)&(0,0)&(0,0)\\ 1&(1,0)&(0,1)&(1,1)\\ 2&(1,1)&(1,0)&(0,1)\\ 3&(0,1)&(1,1)&(1,0)\end{array}$
@@
 Write the binary expansion $n=\sum_{m\geq 0}b_{m}2^{m}$ with $b_{m}\in\{0,1\}$, and define
-$q(n):=\sum_{\begin{subarray}{c}m\geq 0\\ m\ (\mathrm{mod}\ 6)\in\{0,1,3,4\}\end{subarray}}b_{m}2^{m},$ (15)
-so that $q(n)$ is obtained from $n$ by zeroing the bits in positions $m\equiv 2,5\pmod{6}$. Then
+$q(n):=\sum_{\begin{subarray}{c}m\geq 0\\ m\not\equiv 2\ (\mathrm{mod}\ 3)\end{subarray}}b_{m}2^{m},$ (16)
+so that $q(n)$ is obtained from $n$ by zeroing the binary digits in positions $m\equiv 2\pmod{3}$. Then
 $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))\qquad(n\geq 0).$
@@
 Proof.
-In Theorem 34, each base-$4$ digit $d_{k}$ contributes (via $i_{k},j_{k}$) to $\mathcal{M}_{2}(n)$ in a way that depends only on $k\bmod 3$: For $k\equiv 0$, we take $i_{k}\oplus j_{k}$, for $k\equiv 1$ we take only $j_{k}$, and for $k\equiv 2$ we take only $i_{k}$.
-Since $i_{k}$ and $j_{k}$ are precisely the binary digits $b_{2k}$ and $b_{2k+1}$, this means that $\mathcal{M}_{2}(n)$ is the XOR of all binary digits of $n$ except those in positions $2k$ with $k\equiv 1\pmod{3}$ and $2k+1$ with $k\equiv 2\pmod{3}$, i.e., except $m\equiv 2,5\pmod{6}$.
-Thus $\mathcal{M}_{2}(n)$ is the parity of the binary digit-sum of $q(n)$ defined in (15), hence $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$.
-The base-$4$ description of $q(n)$ is the same statement grouped into pairs of bits.
+In Theorem 25, each base-$4$ digit $d_{k}$ contributes to $\mathcal{M}_{2}(n)$ in a way that depends only on $k\bmod 3$: For $k\equiv 0\bmod 3$, we take $i_{k}\oplus j_{k}$, for $k\equiv 1\bmod 3$, we take only $j_{k}$, and for $k\equiv 2\bmod 3$, we take $i_{k}$.
+Since $i_{k}=b_{2k}$ and $j_{k}=b_{2k+1}$, the binary positions $m=2k$ and $m=2k+1$ are kept or omitted, according to the residue class $k\bmod 3$, in the following manner.
+For $k\equiv 0\pmod{3}$, both positions are kept.
+For $k\equiv 1\pmod{3}$, the position $m=2k$ is omitted.
+For $k\equiv 2\pmod{3}$, the position $m=2k+1$ is omitted.
+In each of the cases where a position is omitted, the omitted position $m$ satisfies $m\equiv 2\pmod{3}$.
+Conversely, every value $m$ such that $m\equiv 2\pmod{3}$ falls into one of the two specified cases.
+Thus, the omitted positions are precisely those such that $m\equiv 2\pmod{3}$, and $\mathcal{M}_{2}(n)$ is the parity of the binary digit-sum of $q(n)$ defined in (16), and hence $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$.
+This gives us, in an equivalent way, the desired base-$4$ description of $q(n)$.
 ∎
@@
 Let $p_{a}(n)$ denote the number of distinct length-$n$ factors in the infinite word $a(0)a(1)a(2)\cdots$.
+A length-$n$ factor $u$ of $a$ is called right-special if both $u0$ and $u1$ are factors of $a$.
+We write $\mathrm{RS}_{a}(n)$ in place of the set of right-special factors of $a$ of length $n$.
 Lemma 27.
-For every infinite binary word $a$ and every $n\geq 1$,
-$p_{a}(n+1)-p_{a}(n)=\#\mathrm{RS}_{a}(n),$
-where $\mathrm{RS}_{a}(n)$ denotes the set of length-$n$ factors $u$ of $a$ such that both $u0$ and $u1$ are factors of $a$.
+For every infinite binary word $a$ and every $n\geq 1$, we have that
+$p_{a}(n+1)-p_{a}(n)=\#\operatorname{RS}_{a}(n).$
 Proof.
@@
 ∎
-Lemma 37 (Regularity of the right-special count).
-If $a$ is $k$-automatic, then the function $n\mapsto\#\mathrm{RS}_{a}(n)$ is $k$-regular in the sense of Allouche–Shallit [3], hence effectively computable from any DFAO for $a$.
-When $\mathrm{RS}_{a}(n)$ is additionally bounded, it is $k$-automatic.
+Lemma 28.
+If $a$ is $k$-automatic, then the function $n\mapsto\#\mathrm{RS}_{a}(n)$ is $k$-regular in the sense of Allouche–Shallit [3], and is effectively computable from any DFAO for $a$.
+Moreover, if $\mathrm{RS}_{a}(n)$ is bounded, then it is $k$-automatic.
 Proof.
-The predicate “$u$ is a length-$n$ factor of aa” and the predicate “$u$ is right-special” are first-order definable over $(\mathbb{N},+,V_{k})$ relative to the DFAO for $a$, so the set of encodings of such pairs is regular and recognized by a finite automaton.
-Counting right-special factors of a given length reduces to a finite-automaton computation, giving a $k$-regular sequence.
-For details and background, we again refer to the appropriate texts from Allouche and Shallit [3] and from Shallit [27].
-If the values are bounded, $k$-regularity reduces to $k$-automaticity.
+The predicate “$u$ is a length-$n$ factor of aa” and the predicate “$u$ is right-special” are first-order definable over $(\mathbb{N},+,V_{k})$ relative to a DFAO for $a$, so the set of encodings of such pairs is regular.
+Counting right-special factors of a given length reduces to a finite-automaton computation, giving a $k$-regular sequence [3, Ch.
+16].
+A bounded $k$-regular sequence is $k$-automatic [3, Theorem 16.1.5].
 ∎
-Several claims in this section can be mechanically certified using the Walnut theorem prover [27], which decides first-order properties of automatic sequences in Presburger arithmetic.
-For clarity, and in the hope of developing a deeper understanding of the behavior of meta-automatic sequences, we present “non-Walnut” proofs wherever appropriate.
-5.1 Quarto
-We first record a key property of the Thue–Morse factor language.
+Several statements in this paper can be certified by the Walnut theorem prover [29], which decides first-order properties of automatic sequences in Presburger arithmetic.
+This applies to the defining recurrences and the balancedness of $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ (Sections 3 and 4), and (using Lemma 28) to the right-special factor structure of $\mathcal{Q}$ and $\mathcal{M}_{2}$ in the present section.
+The right-special factor verifications for $\mathcal{M}_{2}$ use the CCL(S) determinization algorithm of Nicol and Frohme [25].
+We give arguments that do not require Walnut wherever this is convenient.
+5.1 Factor complexity of $\mathcal{Q}$
+We first state a property of the factors of $\mathbf{t}$.
 Lemma 29.
-The language $\mathrm{Fac}(\mathbf{t})$ is closed under bitwise complementation, i.e., if $w\in\mathrm{Fac}(\mathbf{t})$, then $\bar{w}\in\mathrm{Fac}(\mathbf{t})$.
+The set $\mathrm{Fac}(\mathbf{t})$ is closed under bitwise complementation: If $w\in\mathrm{Fac}(\mathbf{t})$, then $\bar{w}\in\mathrm{Fac}(\mathbf{t})$.
 Proof.
@@
 Proof.
-By Proposition 16, on each dyadic block $2^{k},2^{k+1})$ the sequence $\mathcal{Q}$ coincides with $\mathbf{t}$ or $1-\mathbf{t}$ according to the parity of $k$.
-A factor of $\mathcal{Q}$ of length $n$ either lies entirely within one block or straddles a dyadic boundary.
-Intra-block factors.
-If the factor lies in an even block it is a factor of $\mathbf{t}$.
-If it lies in an odd block it is a complement of a factor of $\mathbf{t}$, which by Lemma 38 is again a factor of $\mathbf{t}$.
-So all intra-block factors belong to $\mathrm{Fac}(\mathbf{t})$.
-Junction factors.
-A factor straddling boundary $2^{k}$ has the form $u\cdot\bar{v}$ (or $\bar{u}\cdot v$, by a symmetric argument), where $u\cdot v\in\mathrm{Fac}(\mathbf{t})$.
-Such a word need not lie in $\mathrm{Fac}(\mathbf{t})$: If $u\cdot\bar{v}\in\mathrm{Fac}(\mathbf{t})$, then $\bar{u}\cdot v\in\mathrm{Fac}(\mathbf{t})$ by Lemma 38, but this need not hold in general.
-We define $J_{n}$ as the set of those length-$n$ junction factors not already in $\mathrm{Fac}(\mathbf{t})$.
-Since every factor of $\mathbf{t}$ appears in $\mathcal{Q}$ (via even blocks, which contain $\mathbf{t}$ itself), the factor language of $\mathcal{Q}$ is exactly $\mathrm{Fac}(\mathbf{t})\cup J_{n}$, giving $p_{\mathcal{Q}}(n)=p_{\mathbf{t}}(n)+|J_{n}|$.
+By Proposition 8, on each interval $2^{k},2^{k+1})$, the sequence $\mathcal{Q}$ coincides with $\mathbf{t}$ or $1-\mathbf{t}$, according to the parity of $k$.
+The cases give us that a factor of $\mathcal{Q}$ of length $n$ is either a factor of $\mathbf{t}$ or the complement of a factor of t, which, by Lemma 29, is again a factor of $\mathbf{t}$.
+In any case, such factors are in $\mathrm{Fac}(\mathbf{t})$.
+For an interval of the form $2^{k},2^{k+1})$, suppose that the indices associated with a factor of length $n$ properly contains this interval.
+This can only occur when its left endpoint is at a position $2^{k}$ and its right endpoint is at a position $\geq 2^{k+1}$ for some $k$, which forces $n\geq 2^{k}$.
+In every case, such a factor has the form $u_{0}\bar{u}_{1}u_{2}\bar{u}_{3}\cdots$ obtained by concatenating consecutive pieces of $\mathbf{t}$ with alternating bitwise complementation, where the concatenation $u_{0}u_{1}u_{2}u_{3}\cdots$ is a factor of $\mathbf{t}$.
+By Lemma 29, the complementation pattern preserves membership in $\mathrm{Fac}(\mathbf{t})$ for the simplest one-boundary case $u_{0}\bar{u}_{1}$ if and only if $\bar{u}_{0}u_{1}\in\mathrm{Fac}(\mathbf{t})$, and this need not hold in general.
+According to the above definition of $J_{n}$, we have that $J_{n}$ as the set of length-$n$ factors that have at least one index strictly less than $2^{k}$ and at least one index greater than or equal to $2^{k}$ and that are not in $\mathrm{Fac}(\mathbf{t})$.
+Since every factor of $\mathbf{t}$ appears in $\mathcal{Q}$, the set of factors of $\mathcal{Q}$ of length $n$ is exactly $\mathrm{Fac}(\mathbf{t})\cap\{0,1\}^{n}\cup J_{n}$, giving us that $p_{\mathcal{Q}}(n)=p_{\mathbf{t}}(n)+|J_{n}|$.
 ∎
@@
 Proof.
-The value $p_{\mathbf{t}}(2^{k})=3\cdot 2^{k}-2$ is known [6, 23].
-By Proposition 39, $p_{\mathcal{Q}}(2^{k})=p_{\mathbf{t}}(2^{k})+|J_{2^{k}}|$.
-The right-special count $\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n)$ takes values in $\{2,3,4\}$, verified computationally for $2\leq n\leq 3000$ (using 30 00030\,000 terms of $\mathcal{Q}$); a complete proof for all $n$ can be obtained through a Walnut certificate, obtainable from the DFAO for $\mathcal{Q}$.
-It follows from the recurrence
+The value $p_{\mathbf{t}}(2^{k})=3\cdot 2^{k}-2$ is known [6, 24]. By Proposition 30, we have that $p_{\mathcal{Q}}(2^{k})=p_{\mathbf{t}}(2^{k})+|J_{2^{k}}|$. Also, Lemma 27 gives us that
+$\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n).$ (17)
+The right-special count in (17) takes values in $\{2,3,4\}$, and this is certified for all $n\geq 2$ by Walnut. The desired result follows from the recurrence
 $p_{\mathcal{Q}}(2^{k+2})=4\,p_{\mathcal{Q}}(2^{k})+6\qquad(k\geq 4),$
-with $p_{\mathcal{Q}}(16)=50$, which has the unique solution $p_{\mathcal{Q}}(2^{k})=13\cdot 2^{k-2}-2$ and which we verify computationally for $4\leq k\leq 7$. ∎
+with $p_{\mathcal{Q}}(16)=50$, which has the unique solution $p_{\mathcal{Q}}(2^{k})=13\cdot 2^{k-2}-2$, also verified computationally for $4\leq k\leq 7$. ∎
 Remark 32.
-The junction set $J_{2^{k}}$ has size $2^{k-2}=\Theta(2^{k})$: Junction factors account for exactly $\tfrac{1}{4}$ of all length-$2^{k}$ factors of $\mathcal{Q}$ beyond those of $\mathbf{t}$.
-Since $\mathrm{RS}_{\mathcal{Q}}(n)\in\{2,3,4\}$ for all $n\geq 1$, the values of $p_{\mathcal{Q}}$ between dyadic points are completely determined.
-The support of each slope takes a precise regular form.
+The set $J_{2^{k}}$ has size $2^{k-2}$, so factors in $J_{2^{k}}$ account for one quarter of the length-$2^{k}$ factors of $\mathcal{Q}$ beyond those of $\mathbf{t}$.
+Since $\mathrm{RS}_{\mathcal{Q}}(n)\in\{2,3,4\}$ for all $n\geq 1$, the values of $p_{\mathcal{Q}}$ are determined by the support of each slope.
+Lemma 27 is implicit in the following Proposition, according to (18).
 Proposition 33.
-For $m\geq 4$, set $k=\lfloor\log_{2}n\rfloor$ and write $n=2^{k}+i$ with $0\leq i<2^{k}$. Then
+For $n\geq 16$, set $k=\lfloor\log_{2}n\rfloor$ and write $n=2^{k}+i$ with $0\leq i<2^{k}$. Then
 $p_{\mathcal{Q}}(n)=\begin{cases}13\cdot 2^{k-2}-2,&\text{if }i=0;\\ 4n-3\cdot 2^{k-2}-4,&\text{if }1\leq i\leq 2^{k-1};\\ 3n+3\cdot 2^{k-2}-3,&\text{if }2^{k-1}<i\leq 3\cdot 2^{k-2};\\ 2n+5\cdot 2^{k-1}-2,&\text{if }3\cdot 2^{k-2}<i<2^{k}.\end{cases}$
-In other words, on each block $2^{k},2^{k+1})$, the right-special count $\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n)$ equals $4$ for $n\in(2^{k},\,3\cdot 2^{k-1}]$, then $3$ for $n\in(3\cdot 2^{k-1},\,7\cdot 2^{k-2}]$, then $2$ for $n\in(7\cdot 2^{k-2},\,2^{k+1})$.
-The formulas agree at the breakpoints and are continuous across dyadic boundaries.
-In particular, Theorem 40 is the restriction to $i=0$.
+Moreover, we have that
+$\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n).$ (18)
+The left-hand side of (18) equals $4$ for $n\in(2^{k},\,3\cdot 2^{k-1}]$, and $3$ for $n\in(3\cdot 2^{k-1},\,7\cdot 2^{k-2}]$, and $2$ for $n\in(7\cdot 2^{k-2},\,2^{k+1})$.
 Proof.
-A complete proof for all $n$ can be obtained via Walnut, leaving the details to the reader. ∎
-As a consequence of Proposition 42, we have that we have $n+1\leq p_{\mathcal{Q}}(n)=O(n)$ for all $n\geq 1$.
-5.2 Meta(1) and Meta(2)
+This is certified by Walnut. ∎
+By Proposition 33, $n+1\leq p_{\mathcal{Q}}(n)=O(n)$ for all $n\geq 1$.
+5.2 Factor complexity of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$
 Proposition 34.
@@
 Proof.
-The incidence matrices of the morphisms in Corollaries 25 and 32 are
-$M_{\mathcal{M}_{1}}=\begin{pmatrix}1&2&1&0\\ 1&1&1&1\\ 1&1&1&1\\ 1&0&1&2\end{pmatrix},\qquad M_{\mathcal{M}_{2}}=\begin{pmatrix}1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\end{pmatrix}$.
-One checks $M_{\mathcal{M}_{1}}^{2}>0$ and $M_{\mathcal{M}_{2}}>0$ directly, so both morphisms are primitive.
-From the work of Pansiot [24], the fixed point of a primitive $k$-uniform morphism over an alphabet $\mathcal{A}$ has complexity $p(n)=\Theta(n)$.
-Each of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is a coding $\tau(\sigma^{\omega}(a_{0}))$, where $\tau\colon\mathcal{A}\to\{0,1\}$ is non-eventually-constant.
-Since every letter appears with positive frequency and $\tau$ takes both values, distinct factors of $\sigma^{\omega}(a_{0})$ not identified by $\tau$ remain distinct in $\tau(\sigma^{\omega}(a_{0}))$, giving $p_{\tau(\sigma^{\omega})}(n)\geq p_{\sigma^{\omega}}(n)/|\mathcal{A}|^{2}=\Theta(n)$ [3, Proposition 10.9.2].
+The incidence matrices of the morphisms in Corollaries 16 and 23 are
+$M_{\mathcal{M}_{1}}=\begin{pmatrix}1&2&1&0\\ 1&1&1&1\\ 1&1&1&1\\ 1&0&1&2\end{pmatrix},\qquad M_{\mathcal{M}_{2}}=\begin{pmatrix}1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\end{pmatrix}$. (19)
+The incidence matrices in (19) give us that both of the associated morphisms are primitive.
+A result of Pansiot [26] gives us that the fixed point of a primitive $k$-uniform morphism has factor complexity $O(n)$, and the same upper bound holds after applying a coding.
+Neither $\mathcal{M}_{1}$ nor $\mathcal{M}_{2}$ is ultimately periodic, as can be checked from the DFAOs of Theorems 13 and 22, or certified by Walnut.
+By the Morse–Hedlund theorem [3, Theorem 10.2.6], the factor complexity of a non-ultimately-periodic word is at least $n+1$ for every $n\geq 0$.
+Hence both complexities are $\Theta(n)$.
 ∎
-The exact complexity at dyadic lengths, and the complete piecewise formula, are given in the theorems below.
+An explicit evaluation of the factor complexity of $\mathcal{M}_{2}$ is provided below.
 Example 35.
 We obtain the initial values
-$p_{\mathcal{Q}}(1\ldots 15)=(2,4,6,10,12,16,20,23,26,30,34,38,42,45,48),$
+$p_{\mathcal{M}_{1}}(1\ldots 15)=(2,4,6,9,12,17,22,28,34,40,46,52,58,64,70),$
 and
@@
 Proposition 36.
-For all $k\geq 3$,
+For all $k\geq 3$, we have that
 $p_{\mathcal{M}_{2}}(2^{k})=\begin{cases}\dfrac{5}{2}\cdot 2^{k}-2,&\text{if }3\mid k;\\ 3\cdot 2^{k}-2,&\text{otherwise.}\end{cases}$
-In particular, the dyadic slope alternates between $\frac{5}{2}$ and $3$ according to $k\bmod 3$. The three cases modulo $3$ are tied together by the recurrence
-$p_{\mathcal{M}_{2}}(2^{k+3})=8\,p_{\mathcal{M}_{2}}(2^{k})+14\qquad(k\geq 3).$
+Moreover, the recurrence
+$p_{\mathcal{M}_{2}}(2^{k+3})=8\,p_{\mathcal{M}_{2}}(2^{k})+14\qquad(k\geq 3)$
+holds true.
 Proof.
-The formula is certifiable by the Walnut theorem prover, leaving the details to the reader. ∎
-Since $\mathrm{RS}_{\mathcal{M}_{2}}(n)\in\{2,4\}$, the values of $p_{\mathcal{M}_{2}}$ between dyadic points are determined by the support of $\mathrm{RS}_{\mathcal{M}_{2}}=4$.
-This support has a precise structure that extends Theorem 45 to a complete evaluation.
+This is certified by Walnut. ∎
+Since $\mathrm{RS}_{\mathcal{M}_{2}}(n)\in\{2,4\}$, the values of $p_{\mathcal{M}_{2}}$ are determined by the positions at which $\mathrm{RS}_{\mathcal{M}_{2}}(n)=4$.
+The following proposition extends Proposition 36 to a complete evaluation.
 Proposition 37.
-For $n\geq 1$, set $k=\lfloor\log_{2}n\rfloor$ and $r=k\bmod 3$.
-Case $r=2$: On the entire block $n\in[2^{k},2^{k+1}]$,
+We have $p_{\mathcal{M}_{2}}(1)=2$. For $n\geq 2$, set $k=\lfloor\log_{2}n\rfloor$ and $r=k\bmod 3$.
+Case $r=2$: On the interval $2^{k},2^{k+1}]$, we have that
 $p_{\mathcal{M}_{2}}(n)=2n+(2^{k}-2).$
-Case $r=0$: The block $n\in[2^{k},2^{k+1}]$ splits into three linear pieces,
+Case $r=0$: On the interval $2^{k},2^{k+1}]$, we have that
 $p_{\mathcal{M}_{2}}(n)=\begin{cases}2n+(2^{k-1}-2),&\text{if }2^{k}\leq n\leq 2^{k}+1;\\ 4n-(3\cdot 2^{k-1}+4),&\text{if }2^{k}+1\leq n\leq 7\cdot 2^{k-2}+1;\\ 2n+(2^{k+1}-2),&\text{if }7\cdot 2^{k-2}+1\leq n\leq 2^{k+1}.\end{cases}$
-Case $r=1$: The block $n\in[2^{k},2^{k+1}]$ splits into three linear pieces,
+Case $r=1$: On the interval $2^{k},2^{k+1}]$, we have that
 $p_{\mathcal{M}_{2}}(n)=\begin{cases}2n+(2^{k}-2),&\text{if }2^{k}\leq n\leq 2^{k}+1;\\ 4n-(2^{k}+4),&\text{if }2^{k}+1\leq n\leq 3\cdot 2^{k-1}+1;\\ 2n+(2^{k+1}-2),&\text{if }3\cdot 2^{k-1}+1\leq n\leq 2^{k+1}.\end{cases}$
 Proof.
-Again, a proof certificate can be obtained via Walnut, leaving the details to the reader. ∎
-6 Concluding remarks
-A parallel direction replaces the dyadic balancedness given by $a(2n)+a(2n+1)=1$ by the ternary constraint $a(3n)+a(3n+1)+a(3n+2)=2$, while still restricting to binary values in $\{0,1\}$.
-This forces exactly two $1$s and one 0 in each base-$3$ block, giving density $2/3$ on prefixes of length $3^{m}$.
-Meta-automatic selectors of the form a(3n+ra(n))a(3n+r-a(n)) arise naturally, and the $\mathbb{F}_{2}$-linearization mechanism of Theorem 23 extends to this setting: while such selectors are not denestable, the ternary balance constraint may again convert them into fixed-argument $\mathbb{F}_{2}$-affine maps, yielding $3$-automatic sequences.
-A small number of initial configurations appear to produce binary $3$-automatic sequences with small DFAOs; exploring this family and its complexity functions is left for a separate treatment.
+This is certified by Walnut. ∎
+6 Conclusion
+Consider replacing $a(2n)+a(2n+1)=1$ by $a(3n)+a(3n+1)+a(3n+2)=2$, with $a$ taking values in $\{0,1\}$.
+This gives rise to expressions of the form a(3n+ra(n))a(3n+r-a(n)), and the analogue of Lemma 5 converts each such expression into a fixed-argument expression.
+A small number of initial configurations produce $3$-automatic binary sequences with small DFAOs; we plan to treat this family separately.
 7 Appendix
 Minimal DFAO sizes for the meta-automatic sequences considered in this paper are listed in Table 1.
-Table 1: Minimal DFAO size for each specialization. “Sel.” counts the number of meta-automatic selectors among $F$ and $G$. Boldface marks the three sequences studied in this paper.
-$F(n)$ $G(n)$ DFAO Sel. Notes
-$a(n)$ $a(n)$ 22 0 Thue–Morse
-$a(n)$ $a(2n)$ $\mathbf{3}$ $\mathbf{0}$ Quarto $\mathcal{Q}$
-$a(n)$ $a(2n+1)$ 22 0 Thue–Morse
-$a(n)$ $a(2n+a(n))$ 44 11
-$a(n)$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{1}$ Meta(1) $\mathcal{M}_{1}$
-$a(2n)$ $a(n)$ 44 0
+Table 1: Minimal DFAO size for each specialization.
+The column “Nested” counts the number of branch rules ($F$ and $G$) whose right-hand side contains an expression of the form $a(\,\cdot\pm a(n))$.
+Boldface marks the three sequences studied in this paper.
+$F(n)$ $G(n)$ DFAO Nested Notes
+$a(n)$ $a(n)$ 2 0 Thue–Morse
+$a(n)$ $a(2n)$ $\mathbf{3}$ $\mathbf{0}$ $\mathcal{Q}$ (Thue–Morse Quarto)
+$a(n)$ $a(2n+1)$ 2 0 Thue–Morse
+$a(n)$ $a(2n+a(n))$ 4 1
+$a(n)$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{1}$ $\mathcal{M}_{1}$
+$a(2n)$ $a(n)$ 4 0
 $a(2n)$ $a(2n)$ 0 ultimately periodic
-$a(2n)$ $a(2n+1)$ 22 0 Thue–Morse
-$a(2n)$ $a(2n+a(n))$ 33 11
-$a(2n)$ a(2n+1a(n))a(2n+1-a(n)) 33 11
-$a(2n+1)$ $a(n)$ 55 0
-$a(2n+1)$ $a(2n)$ 33 0 complement of $\mathcal{Q}$
-$a(2n+1)$ $a(2n+1)$ 55 0
-$a(2n+1)$ $a(2n+a(n))$ 55 11
-$a(2n+1)$ a(2n+1a(n))a(2n+1-a(n)) 55 11
-$a(2n+a(n))$ $a(n)$ 44 11
-$a(2n+a(n))$ $a(2n)$ 44 11
-$a(2n+a(n))$ $a(2n+1)$ 44 11
-$a(2n+a(n))$ $a(2n+a(n))$ 44 22
-$a(2n+a(n))$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{2}$ Meta(2) $\mathcal{M}_{2}$
-a(2n+1a(n))a(2n+1-a(n)) $a(n)$ 44 11
-a(2n+1a(n))a(2n+1-a(n)) $a(2n)$ 44 11
-a(2n+1a(n))a(2n+1-a(n)) $a(2n+1)$ 44 11
-a(2n+1a(n))a(2n+1-a(n)) $a(2n+a(n))$ 44 22 complement of $\mathcal{M}_{2}$
-a(2n+1a(n))a(2n+1-a(n)) a(2n+1a(n))a(2n+1-a(n)) 44 22
+$a(2n)$ $a(2n+1)$ 2 0 Thue–Morse
+$a(2n)$ $a(2n+a(n))$ 3 1
+$a(2n)$ a(2n+1a(n))a(2n+1-a(n)) 3 1
+$a(2n+1)$ $a(n)$ 5 0
+$a(2n+1)$ $a(2n)$ 3 0 complement of $\mathcal{Q}$
+$a(2n+1)$ $a(2n+1)$ 5 0
+$a(2n+1)$ $a(2n+a(n))$ 5 1
+$a(2n+1)$ a(2n+1a(n))a(2n+1-a(n)) 5 1
+$a(2n+a(n))$ $a(n)$ 4 1
+$a(2n+a(n))$ $a(2n)$ 4 1
+$a(2n+a(n))$ $a(2n+1)$ 4 1
+$a(2n+a(n))$ $a(2n+a(n))$ 4 2
+$a(2n+a(n))$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{2}$ $\mathcal{M}_{2}$
+a(2n+1a(n))a(2n+1-a(n)) $a(n)$ 5 1
+a(2n+1a(n))a(2n+1-a(n)) $a(2n)$ 4 1
+a(2n+1a(n))a(2n+1-a(n)) $a(2n+1)$ 5 1
+a(2n+1a(n))a(2n+1-a(n)) $a(2n+a(n))$ 5 2
+a(2n+1a(n))a(2n+1-a(n)) a(2n+1a(n))a(2n+1-a(n)) 5 2
 8 Acknowledgments
-Claude Opus 4.6 (Anthropic) was used as a computational assistant, in particular for exploring the factor-complexity functions of $\mathcal{Q}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{1}$ of Section 5 and guessing the formulas, and systematically investigating variants of the recursions as in Section 7, which allowed us to select $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ as three sequences of particular interest for our purposes.
+The authors thank the anonymous referee for a careful reading of the manuscript and for helpful suggestions.
+The authors also thank Gandhar Joshi for independently testing the Walnut files.
+Claude (Anthropic) was used as a computational assistant, in particular for exploratory computations, for suggesting candidate formulas later checked by independent scripts, and for checking consistency between the manuscript, the Python script, and the Walnut files.
+The mathematical content, the proofs, and the final wording are the authors’ own.
 The first author thanks Jean-Paul Allouche, Karl Dilcher, and Jeffrey Shallit for useful feedback related to this paper.